### v2

# Finite and full scale localization for the multi-frequency quasi-periodic CMV matrices

Abstract
This paper formulates the finite and full-scale localization for multi-frequency quasi-periodic CMV matrices. This can be viewed as the CMV counterpart to the localization results by Goldstein, Schlag, and Voda [15, 16] on multi-frequency quasi-periodic Schrödinger operators.
keywords:
CMV matrix; Multi-frequency; Anderson localization; No double resonances condition; Factorization of characteristic determinant.
MSC:
[2010] 37A30 , 42C05 , 70G60
journal: aXiv
1 Introduction
Anderson localization (AL) describes the transformation of an impurity doped medium from a conductor to an insulator and this interesting phenomenon was discovered in physical experiments; see the Anderson’s seminal paper [1]. However, AL also holds significant mathematical importance. Mathematically, it implies that the corresponding operator has pure point spectrum with exponentially decaying eigenfunctions. We refer the reader to [11, 22] for the history of AL. Here we would like to mention that AL problems for Schrödinger operators are widely studied and fruitful results have been obtained, especially for quasi-periodic potentials with single frequency; see for example, [2, 4, 5, 8, 11, 12, 23] and references therein.
Given the development of AL theory for Schrödinger operators, it is natural to investigate the Schrödinger operators with multi-frequency quasi-periodic (MF-QP) potentials. But it is a consensus that shifts on a multidimensional torus $\mathbb{T}^{d}$ turned out to be harder to analyze and the theory was less developed. Chulaevsky and Sinai [10] originally established AL for two-frequency Schrödinger operator almost everywhere in phase space. For the frequency of more than two dimensional case, the same result was obtained in [4]. However, these results were established directly in infinite volume, whereas localization in finite volume was not studied. Recently, Goldstein, Schlag and Voda [15, 16] initially developed some analytical techniques and established both the finite and full scale localization of the MF-QP Schrödinger operators. Based on the tools obtained in [15], Goldstein, Schlag and Voda [16] were committed to a class of multi-frequency quasi-periodic Schrödinger operators and they derived a surprising result that the spectrum of this kind of operators is actually a single interval. This proves a conjecture proposed by Chulaevsky and Sinai in [10]. Here we would also like to mention the recent papers [28] and [29] on MF-QP operators.
As for CMV matrices, substantial progress has been made in developing the mathematical theory of localization. In [9], Cedzich and Werner studied the one-dimensional electric quantum walks, they found out that the absolutely continuous spectrum of the quantum walk with a kind of coins is empty. In addition, as a result of the connection between one-dimensional quantum walks and CMV matrices, the AL result for the walk implies AL for the CMV matrices with skew-shift Verblunsky coefficients. In [20], Lin, Guo and Piao extended the results of Bourgain, Goldstein, Schlag [6] and derived the positivity of the Lyapunov exponents and AL for a class of CMV matrices whose Verblunsky coefficients are defined by skew-shift. Wang and Damanik [27] extended the results by Bourgain and Schlag [5] and obtained AL for the quasi-periodic CMV matrices with single frequency. Zhu [30] established both AL and strong dynamical localization for random CMV matrices. Lin, Guo and Piao [21] proved AL for CMV matrices with Verblunsky coefficients generated by hyperbolic toral automorphism. Recently, for CMV matrices with small and quasi-periodic Verblunsky coefficients, Li, Damanik and Zhou [18] proved that the associated spectral measure is purely absolutely continuous on the essential spectrum. In another paper [19], they established Cantor spectrum for some kind of MF-QP CMV matries with Verblunsky coefficients distributed on a circle inside $\partial\mathbb{D}$ by the Gap Labeling Theorem. However, we still do not really understand much about the MF-QP CMV matrices.
To our knowledge, no research paper has addressed the finite and full-scale localization for MF-QP CMV matrices. This paper aims to fill this gap in the literature. To do that, we need large deviation theorem (LDT) for the entries of the transfer matrix. Comparing with the results in [13], there are some difficulties we have to overcome. This is primarily due to the more complex structure of CMV matrices. It is known to all, Schrödinger operator is a real tri-diagonal matrix, while the CMV matrix is a complex five-diagonal matrix, then the analysis about the corresponding characteristic determinant (entries of the transfer matrix) is more complicated. To treat this issue, we decompose the matrix into two parts and consider two cases; see the proof of Lemma 3.15 in detail. Consequently, the large deviation theorem for the entries of the transfer matrix follows.
In addition, considering the restriction of the extended CMV matrix to a finite interval, from [17] and [30], the Poisson formula of CMV matrices depends on the parity of the endpoints of this interval. This also makes the analysis of the covering form of LDT more difficult to some extent. Thus we consider a subinterval of the above finite interval instead, then we get the covering lemma. Since the subinterval is a closed interval, the covering lemma can be applied to the original interval. And the covering form of LDT is obtained.
Furthermore, the restriction of the Schrödinger operator to a finite interval results in a Hermitian matrix with many favorable properties. In contrast, the restriction of the CMV matrix results in a unitary matrix after modifying the boundary conditions, and its properties are less favorable than those of a Hermitian matrix. For example, there is a famous Cauchy Interlacing Theorem for the eigenvalues of Hermitian matrices. But we can not expect such a result for unitary matrices. This directly results in a challenge in estimating the number of zeros for the characteristic determinant within a disk. Thus, the factorization of the characteristic determinant becomes more complex.
In contrast to the results on localization for quasi-periodic CMV matrices with single frequency [27], we provide a more detailed description of the elimination of double resonances. In [27], one just needs to eliminate a set of frequency satisfying double resonances condition by fixing some phase. While in our paper, we find the sets of frequency and phase respectively, then we prove the full scale localization for all frequencies and phases outside these two sets. In contrast to the results on localization for quasi-periodic CMV matrices with single frequency[27], we provide a more detailed description of the elimination of double resonances. Naturally, the process becomes more complex.
The rest of this paper is organized as follows. We describe the settings in Section 2. In Section 3, we give the estimates of the Lyapunov exponent and prove LDT for the transfer matrix and the entries of the transfer matrix respectively. Furthermore, with the aid of Cartan’s estimate and Poisson formula, we get the spectral and covering form of LDT. Based on the results derived in Section 3, we prove the finite scale localization of the MF-QP CMV matrices in Section 4, see Theorem 3.27. We introduce the no double resonances (NDR) condition in Section 5. We study the factorization of the characteristic determinant under NDR condition in Section 6. In Section 7, we derive a result on elimination of double resonances by applying the theory of semialgebraic sets. In Section 8, we study the elimination of double resonances between NDR intervals. Based on the analysis in previous sections, we obtain the full scale localization of the MF-QP CMV matrices in Section 9, see Theorem 6.7. We put some useful definitions and lemmas in the appendix section, that is, Section 10.
2 Preliminaries
In this section, we introduce the preliminaries of orthogonal polynomials on the unit circle (OPUC) that can be seen in detail in [24] and describe the settings we study.
Let $\mathbb{D}=\{z\mid|z|<1\}$ denote the open unit disk in $\mathbb{C}$, and let $\mu$ be a nontrivial probability measure on $\partial\mathbb{D}=\{z\mid|z|=1\}$. Consequently, the functions $1,z,z^{2},\ldots$ are linearly independent in the Hilbert space $L^{2}(\partial\mathbb{D},d\mu)$. Let $\Phi_{n}(z)$ be the monic orthogonal polynomials, that is, $\Phi_{n}(z)=P_{n}[z^{n}]$, $P_{n}\equiv$ projection onto $\{1,z,z^{2},\ldots,z^{n-1}\}^{\bot}$ in $L^{2}(\partial\mathbb{D},d\mu)$. The orthonormal polynomials are then given by
$\varphi_{n}(z)=\frac{\Phi_{n}(z)}{\|\Phi_{n}(z)\|_{\mu}},$
where $\|\cdot\|_{\mu}$ denotes the norm of $L^{2}(\partial\mathbb{D},d\mu)$.
For any polynomial $Q_{n}(z)$ of degree $n$, its Szegő dual $Q_{n}^{*}(z)$ is defined as
$Q_{n}^{*}(z)=z^{n}\overline{Q_{n}(1/\overline{z})}.$
Then the Verblunsky coefficients $\{\alpha_{n}\}_{n=0}^{\infty}$ ($\alpha_{n}\in\mathbb{D}$) obey the equation
$\Phi_{n+1}(z)=z\Phi_{n}(z)-\overline{\alpha}_{n}\Phi_{n}^{*}(z)$ (2.1)
which is known as the Szegő recursion.
Applying the Gram-Schmidt procedure to $1,z,z^{-1},z^{2},z^{-2},\cdots$, we obtain a CMV basis $\{\chi_{j}\}_{j=0}^{\infty}$ of $L^{2}(\partial\mathbb{D},d\mu)$, and the matrix representation of multiplication by $z$ relative to the CMV basis gives rise to the CMV matrix $\mathcal{C}$,
$\mathcal{C}_{ij}=\langle\chi_{i},z\chi_{j}\rangle.$
The matrix $\mathcal{C}$ takes the following form
$\mathcal{C}=\left[\begin{array}[]{ccccccc}\overline{\alpha}_{0}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ \rho_{0}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&\ddots&\ddots&\ddots\end{array}\right],$
where $\rho_{j}=\sqrt{1-|\alpha_{j}|^{2}}$, $j\geq 0$.
The extended CMV matrix is a special five-diagonal doubly-infinite matrix in the standard basis of $L^{2}(\partial\mathbb{D},d\mu)$ as described in [24, Subsection 4.5] and [25, Subsection 10.5], and is given by
$\mathcal{E}=\left[\begin{array}[]{cccccccc}\ddots&\ddots&\ddots&&&&&\\ &-\overline{\alpha}_{0}\alpha_{-1}&\overline{\alpha}_{1}\rho_{0}&\rho_{1}\rho_{0}&&&&\\ &-\rho_{0}\alpha_{-1}&-\overline{\alpha}_{1}\alpha_{0}&-\rho_{1}\alpha_{0}&&&&\\ &&\overline{\alpha}_{2}\rho_{1}&-\overline{\alpha}_{2}\alpha_{1}&\overline{\alpha}_{3}\rho_{2}&\rho_{3}\rho_{2}&&\\ &&\rho_{2}\rho_{1}&-\rho_{2}\alpha_{1}&-\overline{\alpha}_{3}\alpha_{2}&-\rho_{3}\alpha_{2}&&\\ &&&&\overline{\alpha}_{4}\rho_{3}&-\overline{\alpha}_{4}\alpha_{3}&\overline{\alpha}_{5}\rho_{4}&\\ &&&&\rho_{4}\rho_{3}&-\rho_{4}\alpha_{3}&-\overline{\alpha}_{5}\alpha_{4}&\\ &&&&&\ddots&\ddots&\ddots\end{array}\right].$
In this paper, we consider a sequence of Verblunsky coefficients generated by an analytic function $\alpha(\cdot):\mathbb{T}^{d}\rightarrow\mathbb{D}$, $\alpha_{n}(x)=\alpha(T^{n}x)=\alpha(x+n\omega)$, $d>1$. Here, $T$ is the shift map defined by $Tx=x+\omega$, $\mathbb{T}:=\mathbb{R}/\mathbb{Z}$, and $x,\omega\in\mathbb{T}^{d}$ are referred to as phase and frequency respectively. Moreover, the frequency vector $\omega\in\mathbb{T}^{d}$ satisfies the standard Diophantine condition
$\|k\cdot\omega\|\geq\frac{p}{|k|^{q}}$ (2.2)
[…]
Assume that the sampling function $\alpha(x)$ satisfies
[…]
and can be extend complex analytically to
$\mathbb{T}^{d}_{h}:=\{x+iy:x\in\mathbb{T}^{d},y\in\mathbb{R}^{d},|y|<h\}$
with some $h>0$. Define its norm as
$\|\alpha\|_{h}=\sup_{|y|<h}|\alpha(x+iy)|,$
and let
$C_{\alpha}=\log\frac{2}{\sqrt{1-\|\alpha\|^{2}_{h}}}.$ (2.4)
The Szegő recursion is then equivalent to
$\rho_{n}(x)\varphi_{n+1}(z)=z\varphi_{n}(z)-\overline{\alpha}_{n}(x)\varphi_{n}^{*}(z),$ (2.5)
where $\rho_{n}(x)=\rho(x+n\omega)$ and $\rho(x)=(1-|\alpha(x)|^{2})^{1/2}$.
Applying the Szegő dual to both sides of (2.5), we obtain
$\rho_{n}(x)\varphi_{n+1}^{*}(z)=\varphi_{n}^{*}(z)-\alpha_{n}(x)z\varphi_{n}(z).$ (2.6)
Equations (2.5) and (2.6) can be can be expressed as
$\left[\begin{array}[]{c}\varphi_{n+1}\\ \varphi_{n+1}^{*}\end{array}\right]=S(\omega,z;x+n\omega)\left[\begin{array}[]{c}\varphi_{n}\\ \varphi_{n}^{*}\end{array}\right],$
where
$S(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}z&-\overline{\alpha}(x)\\ -\alpha(x)z&1\end{array}\right].$
Since $\det S(\omega,z;x)=z$,it is preferable to study the determinant $1$ matrix
$M(\omega,z;x)=\frac{1}{\rho(x)}\left[\begin{array}[]{cc}\sqrt{z}&-\frac{\overline{\alpha}(x)}{\sqrt{z}}\\ -\alpha(x)\sqrt{z}&\frac{1}{\sqrt{z}}\end{array}\right]\in\mathbb{S}\mathbb{U}(1,1),$
known as the Szegő cocycle map.
The monodromy matrix (or $n$-step transfer matrix) is then defined by
$M_{n}(\omega,z;x)=\prod_{j=n-1}^{0}M(\omega,z;x+j\omega).$ (2.7)
From (2.7), it follows directly that
$M_{n_{1}+n_{2}}(\omega,z;x)=M_{n_{2}}(\omega,z;x+n_{1}\omega)M_{n_{1}}(\omega,z;x)$
and
$\log\|M_{n_{1}+n_{2}}(\omega,z;x)\|\leq\log\|M_{n_{1}}(\omega,z;x)\|+\log\|M_{n_{2}}(\omega,z;x+n_{1}\omega)\|.$ (2.8)
Define
$u_{n}(\omega,z;x):=\frac{1}{n}\log\|M_{n}(\omega,z;x)\|$ (2.9)
and
[…]
Integrating the inequality (2.8) with respect to $x$ over $\mathbb{T}^{d}$ yields
$L_{n_{1}+n_{2}}(\omega,z)\leq\frac{n_{1}}{n_{1}+n_{2}}L_{n_{1}}(\omega,z)+\frac{n_{2}}{n_{1}+n_{2}}L_{n_{2}}(\omega,z).$
This implies that
$L_{n}(\omega,z)\leq L_{m}(\omega,z)\ \ {\rm if}\ \ m<n,\;m|n$
and
$L_{n}(\omega,z)\leq L_{m}(\omega,z)+C\frac{m}{n}\ \ {\rm if}\ \ m<n.$
For any irrational $\omega\in\mathbb{T}^{d}$, the transformation $x\rightarrow x+\omega$ is ergodic. Notice that (2.8) implies that $\log\|M_{n}(\omega,z;x)\|$ is subadditive, Kingman’s subadditive ergodic theorem guarantees the existence of the limit
$L(\omega,z)=\lim_{n\rightarrow\infty}L_{n}(\omega,z),$ (2.11)
known as the Lyapunov exponent. Throughout this paper, $\gamma>0$ denotes the lower bound of the Lyapunov exponent.
Moreover, the Furstenberg-Kesten theorem states that the limit
$\lim_{n\rightarrow\infty}u_{n}(\omega,z;x)=\lim_{n\rightarrow\infty}L_{n}(\omega,z)=L(\omega,z).$ (2.12)
exists for $a.e.\;x$. Note that
$0\leq\log\|M_{n}(\omega,z;x)\|\leq C(\alpha_{i},z)n$ (2.13)
and hence
$0\leq L_{n}(\omega,z)\leq C(\alpha_{i},z)n$ (2.14)
for $i\in\mathbb{Z}$.
3 Some Useful Lemmas
3.1 LDT Estimate for the Monodromy
To establish LDT, we rely on the following lemma.
Lemma 3.1.
[14, Proposition 9.1] Let $d$ be a positive integer. Suppose $u:D(0,2)^{d}\rightarrow[-1,1]$ is subharmonic in each variable; i.e., $z\rightarrow u(z_{1},u_{2},\ldots,z_{d})$ is subharmonic for any choice of $(z_{2},\ldots,z_{d})\in D(0,2)^{d-1}$ and similarly for each of the other variables. Assume furthermore that for some $n\geq 1$
$\sup_{\theta\in\mathbb{T}^{d}}|u(\theta+\omega)-u(\theta)|<\frac{1}{n}.$ (3.1)
Then there exist $\sigma>0$, $\tau>0$, and $c_{0}$ only depending on $d$ and $\varepsilon_{1}$ such that
$\mathrm{mes}\{\theta\in\mathbb{T}^{d}:|u(\theta)-\langle u\rangle|>n^{-\tau}\}<\exp(-c_{0}n^{\sigma}).$ (3.2)
[…]
Based on the above lemma, obtain the following LDT, a fundamental tool in localization theory.
Theorem 3.2.
Suppose $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. There exist constants $\sigma=\sigma(p,q)$, $\tau=\tau(p,q)$ with $\sigma,\tau\in(0,1)$, and $C_{0}=C_{0}(p,q,h)$ such that for $n\geq 1$,
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x)\|-nL_{n}(\omega,z)|>n^{1-\tau}\}<\exp(-C_{0}n^{\sigma}).$ (3.3)
Proof.
Fix a dimension $d$ and $z$. For any $x\in D(0,2)^{d}$, define
$u_{n}(x)=\frac{1}{n}\log\|M_{n}(\omega,z;x)\|.$
Then $u_{n}$ is a continuous subharmonic function bounded by $1$ in $D(0,1)^{d}$. Moreover, $u_{n}$ satisfies the conditions in the above lemma, and (3.3) follows directly from (3.2). ∎
3.2 Estimation of the Lyapunov Exponent
Lemma 3.3.
Assume $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. For any $n\geq 2$, we have
$0\leq L_{n}(\omega,z)-L(\omega,z)<C\frac{(\log n)^{1/\sigma}}{n},$
where $C=C(p,q,z,\gamma)$ and $\sigma$ is as in LDT.
Proof.
Clearly, $0\leq L_{n}(\omega,z)\leq C(z)$ for all $n$. Let $k$ be a positive integer satisfying $k\gamma>16C(z)$. For $n>10$, set $l=[C_{1}(\log n)^{1/\sigma}]$ with some large $C_{1}$. Consider the integers $l_{0},2l_{0},\ldots,2^{k}l_{0}$. There exists some $0\leq j<k$ such that (with $l_{j}=2^{j}l_{0}$)
$L_{l_{j}}(z)-L_{l_{j+1}}(z)<\frac{\gamma}{16}.$ (3.4)
Otherwise, then $C(z)>L_{l_{0}}(z)-L_{2^{k}l_{0}}(z)\geq\frac{k\gamma}{16}>C(z)$, leading to a contradiction. Set $l=2^{j}l_{0}$ where $j$ satisfies (3.4).
It is straightforward to verify that to verify that $M(\omega,z;x)$ is conjugate to an $SL(2,\mathbb{R})$ matrix $T(\omega,z;x)$ via Q i.e.,
$Q=-\frac{1}{1+i}\left[\begin{array}[]{cc}1&-i\\ 1&i\end{array}\right]\in\mathbb{U}(2),$
that is,
$T(\omega,z;x)=Q^{*}M(\omega,z;x)Q\in SL(2,\mathbb{R}),$
where “*” denotes the conjugate transpose of a matrix.
We now apply the Avalanche Principle (AP) (see Appendix) to the matrices $B_{j}=T_{l}(\omega,z;x+(j-1)l\omega)$ for $1\leq j\leq m=[n/l]$, where $\mu=\exp(l\gamma/2)$. Note that $\mu>n^{2}$ if $C_{1}>4/\gamma$. By LDT, the phases $x$ satisfying
$\min_{1\leq j\leq m}\|B_{j}\|\geq\exp(lL_{l}(\omega,z)-l^{1-\tau})>\exp(l\gamma/2)=\mu>n^{2}$
form a set $\mathcal{G}_{1}$ with measure at most $m\exp(-C_{0}l^{\sigma})\leq n\exp(-C_{0}(2^{j}l_{0})^{\sigma})<n^{-2}$, provided $C_{1}>n^{1/\sigma}$.
Moreover, combining (3.4) and LDT, we have
$\max_{1\leq j<m}\Big{|}\log\|B_{j+1}\|+\log\|B_{j}\|-\log\|B_{j+1}B_{j}\|\Big{|}$ $\leq 2l(L_{l}(z)+\frac{\gamma}{32})-2l(L_{2l}(z)-\frac{\gamma}{32})$
$=2l[(L_{l}(z)-L_{2l}(z))+\frac{\gamma}{16})]$
$\leq\exp(l\gamma/4)=\frac{1}{2}\log\mu$
up to a set $\mathcal{G}_{2}$ of $x$ with measure at most $2m\exp(-C_{0}l^{\sigma})\leq 2n\exp(-C_{0}(l)^{\sigma})<n^{-2}$, provided $C_{1}>n^{1/\sigma}$. Let $\mathcal{G}=\mathcal{G}_{1}\cup\mathcal{G}_{2}$. Then $\mathrm{mes}\mathcal{G}<2n^{-2}$, and (7.1), (7.2) hold for all $x\in\mathcal{G}$.
By (AP), we have
$\Big{|}\log\|B_{m}\cdots B_{1}\|+\sum_{j=2}^{m-1}\log\|B_{j}\|-\sum_{j=1}^{m-1}\log\|B_{j+1}B_{j}\|\Big{|}<C\frac{n}{\mu}<Cn^{-1}$
for all $x\in\mathcal{G}$. That is,
$\Big{|}\log\|T_{lm}(\omega,z;x)+\sum_{j=2}^{m-1}\log\|T_{l}(\omega,z;x+jl\omega)\|$
$\ \ \ -\sum_{j=1}^{m-1}\log\|T_{l}(\omega,z;x+(j+1)l\omega)T_{l}(\omega,z;x+jl\omega)\|\Big{|}<Cn^{-1}$
for all $x\in\mathcal{G}$.
Similarly, for $\log\|T_{lm}(\omega,z;x+lm\omega)\|$ and $\log\|T_{lm}(\omega,z;x)\|$, the phases $x$ satisfying
$\Big{|}\|\log\|T_{2lm}(\omega,z;x)\|-\|\log\|T_{lm}(\omega,z;x+lm\omega)\|-\|\log\|T_{lm}(\omega,z;x)\|$
$+\|\log\|T_{l}(\omega,z;x+lm\omega)\|+\|\log\|T_{l}(\omega,z;x+(m-1)l\omega)\|$
$-\|\log\|T_{l}(\omega,z;x+lm\omega)T_{l}(\omega,z;x+(m-1)l\omega)\|\Big{|}\leq\frac{C}{n}$ (3.5)
form a set with measure at most $Cn^{-2}$.
Since
$\big{|}\|\log\|T_{n}(\omega,z;x)\|-\log\|T_{lm}(\omega,z;x)\|\big{|}\leq C(\alpha_{n},\alpha_{lm},z)n$ (3.6)
and
$\big{|}\|\log\|T_{l}(\omega,z;x)\|\big{|}\leq C(z)l,$ (3.7)
we can conclude from (3.5) that
$\big{|}\|\log\|T_{2n}(\omega,z;x)\|-\log\|T_{n}(\omega,z;x+n\omega)\|-\log\|T_{n}(\omega,z;x)\|\big{|}\leq C(\log n)^{1/\sigma}$
up to a set of $x$ not exceeding $Cn^{-2}$ in measure.
Integrating the above inequality over $x$, then we have
$|L_{2n}(z)-L_{n}(z)|\leq C\frac{(\log n)^{1/\sigma}}{n},$
where $C=C(\gamma,z,\omega)$.
We can finally prove this lemma by summing over $2^{k}n$ ∎
Lemma 3.4.
Let $n\geq 1$, $(\omega_{i},z_{i},x_{i})\in\mathbb{C}^{d}\times\partial\mathbb{D}\times\mathbb{C}^{d}$, $i=1,2$, such that
$|\mathrm{Im}x_{i}|<h,\quad n|\mathrm{Im}\omega_{i}|<h.$
Then
$\|M_{n}(\omega_{1},z_{1};x_{1})-M_{n}(\omega_{2},z_{2};x_{2})\|\leq\Big{(}\frac{4}{\sqrt{1-\|\alpha\|}_{h}}\Big{)}^{n}(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|).$
In particular, we have
$\big{|}\log\|M_{n}(\omega_{1},z_{1};x_{1})\|-\log\|M_{n}(\omega_{2},z_{2};x_{2})\|\big{|}\leq\Big{(}\frac{4}{\sqrt{1-\|\alpha\|}_{h}}\Big{)}^{n}(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|),$
provided the right-hand side is less than $\frac{1}{2}$.
Proof.
Let
$M_{k,n}=\frac{1}{\rho(x_{k}+n\omega_{k})}\left[\begin{array}[]{cc}\sqrt{z_{k}}&-\frac{\overline{\alpha}(x_{k}+n\omega_{k})}{\sqrt{z_{k}}}\\ -\alpha(x_{k}+n\omega_{k})&1/\sqrt{z_{k}}\end{array}\right],\quad k=1,2.$
Then
$M_{n}(\omega_{1},z_{1};x_{1})-M_{n}(\omega_{2},z_{2};x_{2})$ $=\sum_{j=1}^{n}M_{2,n}\cdots M_{2,j+1}(M_{1,j}-M_{2,j})M_{1,j-1}\cdots M_{1,1}$
$=\sum_{j=1}^{n}M_{n-j}(\omega_{2},z_{2};x_{2})(M_{1,j}-M_{2,j})M_{j-1}(\omega_{1},z_{1};x_{1}).$
It’s not difficult to check that $M_{k,n}$ is conjugate to an $SL(2,\mathbb{R})$ matrix $T_{k,n}$ through $Q$, that is,
$T_{k,n}=Q^{*}M_{k,n}Q.$
According to the polar decomposition of $SL(2,\mathbb{C})$ matrices,
$\mathrm{tr}(T_{k,n}^{*}T_{k,n})=|a|^{2}+|b|^{2}+|c|^{2}+|d|^{2}=\|T_{k,n}\|^{2}+\|T_{k,n}\|^{-2},$
where $a,b,c,d$ are the entries of matrix $T_{k,n}$. Then a direct calculation implies
$\|T_{k,n}\|+\|T_{k,n}\|^{-1}$ $=\|M_{k,n}\|+\|M_{k,n}\|^{-1}$
$=\sqrt{\mathrm{tr}(M_{k,n}^{*}M_{k,n})+2}$
$=\frac{\sqrt{2+2|\alpha(x_{k}+n\omega_{k})|^{2}+2\rho^{2}(x_{k}+n\omega_{k})}}{\rho(x_{k}+n\omega_{k})}$
$=\frac{2}{\sqrt{1-|\alpha(x_{k}+n\omega_{k})|^{2}}}.$
Thus, $\sup\limits_{|y|<h}\|T_{k,n}(\omega_{k},z_{k};x_{k}+iy_{k})\|$, $\sup\limits_{|y|<h}\|T_{k,n}(\omega_{k},z_{k};x_{k}+iy_{k})\|^{-1}\leq\frac{2}{\sqrt{1-\|\alpha\|_{h}^{2}}}$ and
$\sup\limits_{|y|<h}\frac{1}{n}\log\|T_{n}(\omega,z;x)\|=\sup\limits_{|y|<h}\frac{1}{n}\log\|M_{n}(\omega,z;x)\|\leq\log\frac{2}{\sqrt{1-\|\alpha\|_{h}^{2}}}=C_{\alpha},$
where $T_{n}(\omega,z;x)=\mathop{\prod}\limits_{j=n-1}^{0}T(\omega,z;x+j\omega)$, $T(\omega,z;x+j\omega)=Q^{*}M(\omega,z;x+j\omega)Q$.
Then according to the Mean Value Theorem, one can obtain that
$\|M_{n}(\omega_{1},z_{1};x_{1})-M_{n}(\omega_{2},z_{2};x_{2})\|\leq\Big{(}\frac{4}{\sqrt{1-\|\alpha\|^{2}}_{h}}\Big{)}^{n}(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|).$
Due to the fact that $|\log x|\leq 2|x-1|$ provided $|x-1|\leq\frac{1}{2}$, we get
$\big{|}\log\|M_{n}(\omega_{1},z_{1};x_{1})\|-\log\|M_{n}(\omega_{2},z_{2};x_{2})\|\big{|}$ $\leq\Big{|}\frac{\|M_{n}(\omega_{1},z_{1};x_{1})\|}{\|M_{n}(\omega_{2},z_{2};x_{2})\|}-1\Big{|}$
$\leq\frac{\|M_{n}(\omega_{1},z_{1};x_{1})-M_{n}(\omega_{2},z_{2};x_{2})\|}{\|M_{n}(\omega_{2},z_{2};x_{2})\|}$
provided the right-hand side is less than $\frac{1}{2}$.
Since $\|M_{n}(\omega,z;x)\|\geq 1$ for any $\omega,z,x$, then we have
$\big{|}\log\|M_{n}(\omega_{1},z_{1};x_{1})\|-\log\|M_{n}(\omega_{2},z_{2};x_{2})\|\big{|}$
$\leq\|M_{n}(\omega_{1},z_{1};x_{1})-M_{n}(\omega_{2},z_{2};x_{2})\|$
$\leq\Big{(}\frac{4}{\sqrt{1-\|\alpha\|^{2}}_{h}}\Big{)}^{n}(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|).$
Based on (AP), we can refine the above lemma.
Lemma 3.5.
Suppose $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$, and $L(\omega_{0},z_{0})>\gamma>0$. Let $\sigma$ be as in LDT and let $A$ be a constant such that $\sigma A\geq 1$. Then there exists a set $\mathcal{B}_{n,\omega_{0},z_{0}}$, $\mathrm{mes}(\mathcal{B}_{n,\omega_{0},z_{0}})<\exp(-C_{0}(\log n)^{\sigma A})$, such that the following holds. For all $n\geq N_{0}(\alpha_{i},p,q,\gamma,A)$,
$\big{|}\log\|M_{n}(\omega,z;x)\|-\log\|M_{n}(\omega_{0},z_{0};x_{0})\|\big{|}<\exp(-(\log n)^{A})$ (3.8)
for any $x_{0}\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega_{0},z_{0}}$ and $(\omega,z,x)\in\mathbb{C}^{d}\times\partial\mathbb{D}\times\mathbb{C}^{d}$ such that
$|\omega-\omega_{0}|,\,|z-z_{0}|,\,|x-x_{0}|<\exp(-(\log n)^{4A}).$ (3.9)
Proof.
Let $l\simeq(\log n)^{A+1}$, $m=[n/l]$, $B_{j}=T_{l}(\omega,z;x+(j-1)l\omega)$ be as in the proof of Lemma 3.3. For convenience, we also assume that $n=ml$.
Take $\mathcal{B}_{n,\omega_{0},z_{0}}\subset\mathbb{T}^{d}$ be the set in (3.3). By LDT and Lemma 3.3, we get that
$\min_{1\leq j\leq m}\|B_{j}(\omega_{0},z_{0};x_{0})\|\geq\exp(lL_{l}(\omega_{0},z_{0})-l^{1-\tau})\geq\exp(l\gamma/2)>m,$
$\max_{1\leq j<m}(\log\|B_{j+1}(\omega_{0},z_{0};x_{0})\|+\log\|B_{j}(\omega_{0},z_{0};x_{0})\|-\log\|B_{j+1}(\omega_{0},z_{0};x_{0})B_{j}(\omega_{0},z_{0};x_{0})\|)\leq\frac{l\gamma}{16}$
for all $x_{0}$ outside $\mathcal{B}_{n,\omega_{0},z_{0}}$ with
$\mathrm{mes}(\mathcal{B}_{n,\omega_{0},z_{0}})\lesssim m\exp(-C_{0}l^{\sigma})<\exp(-C_{0}(\log n)^{\sigma A}),$
provided that $\sigma A\geq 1$. Take $x_{0}\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega_{0},z_{0}}$. Applying (AP) with $\mu=\exp(l\gamma/2)$, we have
$\Big{|}\log\|M_{m}(\omega_{0},z_{0};x_{0})\|+\sum_{j=2}^{m-1}\log\|B_{j}(\omega_{0},z_{0};x_{0})\|-\sum_{j=1}^{m-1}\log\|B_{j+1}(\omega_{0},z_{0};x_{0})B_{j}(\omega_{0},z_{0};x_{0})\|\Big{|}$
$<Cn\exp(-l\gamma/2).$ (3.10)
Take $\omega,z,x$ satisfying (3.9). From Lemma 3.4, we know that
$\min_{1\leq j\leq m}\|B_{j}(\omega,z;x)\|\geq\exp(l\gamma/4)>m$
and
$\max_{1\leq j<m}(\log\|B_{j+1}(\omega,z;x)\|+\log\|B_{j}(\omega,z;x)\|-\log\|B_{j+1}(\omega,z;x)B_{j}(\omega,z;x)\|)\leq\frac{l\gamma}{8}.$
Now applying (AP) again with $\mu=\exp(l\gamma/4)$, we have
$\Big{|}\log\|M_{m}(\omega,z;x)\|+\sum_{j=2}^{m-1}\log\|B_{j}(\omega,z;x)\|-\sum_{j=1}^{m-1}\log\|B_{j+1}(\omega,z;x)B_{j}(\omega,z;x)\|\Big{|}$
$<Cn\exp(-l\gamma/4).$ (3.11)
Subtracting (3.2) from (3.2) and applying Lemma 3.4, we get
$\big{|}\log\|M_{m}(\omega,z;x)\|-\log\|M_{m}(\omega_{0},z_{0};x_{0})\|\big{|}$ $\leq\Big{(}\frac{4}{\sqrt{1-\|\alpha\|}_{h}}\Big{)}^{n}(|\omega-\omega_{0}|+|z-z_{0}|+|x-x_{0}|)$
$\leq 3\Big{(}\frac{4}{\sqrt{1-\|\alpha\|}_{h}}\Big{)}^{n}\exp(-(\log n)^{4A})$
$<\exp(-(\log n)^{A}).$
Lemma 3.6.
Suppose $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$, and $L(\omega_{0},z_{0})>\gamma>0$. Let $\sigma$ be as in LDT and $A$ be a constant satisfying $\sigma A\geq 1$. Then for all $n\geq N_{0}(p,q,\gamma,A)$,
$|L_{n}(\omega,z)-L_{n}(\omega_{0},z_{0})|<\exp(-(\log n)^{\sigma A})$
provided that $|\omega-\omega_{0}|$, $|z-z_{0}|<\exp(-(\log n)^{4A})$.
Proof.
Due to
$|L_{n}(\omega,z)-L_{n}(\omega_{0},z_{0})|$ $=\Big{|}\frac{1}{n}\Big{(}\int_{\mathbb{T}^{d}}\log\|M_{n}(\omega,z;x)\|dx-\int_{\mathbb{T}^{d}}\log\|M_{n}(\omega_{0},z_{0};x_{0})\|dx\Big{)}\Big{|}$
[…]
then according to Lemma 3.5 we have
$|L_{n}(\omega,z)-L_{n}(\omega_{0},z_{0})|<\exp(-(\log n)^{\sigma A}).$
Lemma 3.7.
Assume $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. Let $\sigma$ be as in LDT. There exists $\varepsilon_{0}=\varepsilon_{0}(p,q,z_{0},\gamma)$ such that if $|z-z_{0}|<\varepsilon_{0}$, then $L(\omega_{0},z)>\frac{\gamma}{2}$ and
$|L(\omega_{0},z)-L(\omega_{0},z_{0})|\leq\exp(\frac{1}{2}(-\log|z-z_{0}|)^{\sigma/4}).$
Furthermore, if $\omega\in\mathbb{T}^{d}(p,q)\cap(\omega_{0}-\varepsilon_{0},\omega_{0}+\varepsilon_{0})$, then $L(\omega,z_{0})>\frac{\gamma}{2}$ and
$|L(\omega,z_{0})-L(\omega_{0},z_{0})|\leq\exp(\frac{1}{2}(-\log|\omega-\omega_{0}|)^{\sigma/4}).$
Proof.
Combining Lemma 3.3 and Lemma 3.6, we have that
$|L(\omega_{0},z)-L(\omega_{0},z_{0})|$ $<C\frac{(\log n)^{1/\sigma}}{n}\leq\exp(-\frac{1}{2}\log(n+1))$
$\leq\exp(\frac{1}{2}(-\log|z-z_{0}|)^{\sigma/4})<\frac{\gamma}{2},$
provided
$\exp(-(\log(n+1))^{4/\sigma})\leq|z-z_{0}|\leq\exp(-(\log n)^{4/\sigma})$
for $n\geq n_{0}(p,q,z_{0},\gamma)$. Then we get the first statement by taking $\varepsilon_{0}=\exp(-(\log n_{0})4/\sigma)$. Similarly, we can obtain the second conclusion. ∎
Lemma 3.8.
[13, Lemma 4.1] Let $10<\rho<$ 1 and suppose $f$ is subharmonic on
$\mathcal{A}_{\rho}:=\{x+iy:x\in\mathbb{T},|y|<\rho\},$
[…]
Corollary 3.9.
Let $\omega\in\mathbb{T}^{d}$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. There exists $C=C(z)$ such that
$|L_{n}(\omega,z;y)-L_{n}(\omega,z)|\leq C\sum_{i=1}^{d}|y_{i}|$
for all $|y|<h$ uniformly in $n$. Particularly, the bound holds equally true with $L$ in place of $L_{n}$.
Proof.
When $d=1$, let $f(x+iy)=\frac{1}{C(z)n}\log\|M_{n}(\omega,z;y)\|$. According to the above lemma,
$|L_{n}(\omega,z;y)-L_{n}(\omega,z)|$
$=\Big{|}\int_{\mathbb{T}}\frac{1}{n}\log\|M_{n}(\omega,z;x+iy)\|dx-\int_{\mathbb{T}}\frac{1}{n}\log\|M_{n}(\omega,z;x)\|dx\Big{|}$
$=C(z)\Big{|}\int_{\mathbb{T}}\frac{1}{C(z)n}\log\|M_{n}(\omega,z;x+iy)\|dx-\int_{\mathbb{T}}\frac{1}{C(z)n}\log\|M_{n}(\omega,z;x)\|dx\Big{|}$
$\leq C(z)C_{h}|y|.$
Now we check the statement for $d=2$. Let
$f(x_{1}+iy_{1};y_{2}):=\int_{\mathbb{T}}\frac{1}{C(z)n}\log\|M_{n}(\omega,z;x_{1}+iy_{1},x_{2}+iy_{2})\|dx_{2}.$
By Lemma 3.8,
$|f(x_{1}+iy_{1};y_{2})-f(x_{1}+iy_{1};0)|\leq C|y_{2}|.$
Fix $y_{2}$. $f(x_{1}+iy_{1};y_{2})$ is a subharmonic function. Lemma 3.8 implies
[…]
Thus,
$|L_{n}(\omega,z;y)-L_{n}(\omega,z)|$
$=\Big{|}\int_{\mathbb{T}^{2}}\frac{1}{n}\log\|M_{n}(\omega,z;x+iy)\|dx-\int_{\mathbb{T}^{2}}\frac{1}{n}\log\|M_{n}(\omega,z;x)\|dx\Big{|}$
$=C(z)\Big{|}\int_{\mathbb{T}}\int_{\mathbb{T}}\frac{1}{C(z)n}\log\|M_{n}(\omega,z;x_{1}+iy_{1},x_{2}+iy_{2})\|dx_{2}dx_{1}$
$\quad-\int_{\mathbb{T}}\int_{\mathbb{T}}\frac{1}{C(z)n}\log\|M_{n}(\omega,z;x_{1},x_{2})\|dx_{2}dx_{1}\Big{|}$
[…]
$\leq C(|y_{1}+y_{2}|).$
For general $d$, the proof is similar with the aid of induction over $d$. ∎
Lemma 3.10.
Assume $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$, and $L(\omega,z)>\gamma>0$. Then for all $n\geq 1$,
$\sup\limits_{x\in\mathbb{T}^{d}}\log\|M_{n}(\omega,z;x)\|\leq nL_{n}(\omega,z)+Cn^{1-\tau},$ (3.12)
where $C=C(p,q,z,\gamma)$ and $\tau$ as in LDT.
Proof.
To check (3.12), we only need to consider the case that $n$ is large enough because for smaller $n$ we can take $C$ large enough.
According to LDT,
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x+iy)\|-nL_{n}(\omega,z;y)|>n^{1-\tau}\}<\exp(-C_{0}n^{\sigma}).$ (3.13)
By Corollary 3.9, we can get
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x+iy)\|-nL_{n}(\omega,z)|>2n^{1-\tau}\}<\exp(-C_{0}n^{\sigma})$ (3.14)
provided $|y|<\frac{1}{n}$. With the help of the sub-mean value property applicable to subharmonic functions, we obtain that
$\log\|M_{n}(\omega,z;x)\|\leq(\pi r^{2})^{-d}\int_{\mathcal{P}}\log\|M_{n}(\omega,z;\xi+iy)\|d\xi dy,$ (3.15)
where the polydisk $\mathcal{P}=\mathop{\prod}\limits_{j=1}^{d}D(x_{j},r)$, $x=(x_{1},\ldots,x_{d})$, $r=\frac{1}{n}$.
Let $\mathcal{A}_{y}\subset\mathbb{T}^{d}$ be the set in (3.14), i.e.,
$\mathcal{A}_{y}=\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x+iy)\|-nL_{n}(\omega,z)|>2n^{1-\tau}\}.$
Let $\mathcal{A}=\{(\xi,y)\in[0,1]^{d}\times(-r,r)^{d}:\xi\in\mathcal{A}_{y}\}$. From (3.14), we have
$(\pi r^{2})^{-d}\int_{\mathcal{P}\backslash\mathcal{A}}\log\|M_{n}(\omega,z;\xi+iy)\|d\xi dy\leq nL_{n}(\omega,z)+2n^{1-\tau}.$ (3.16)
Due to (2.13),
$(\pi r^{2})^{-d}\int_{\mathcal{P}\cap\mathcal{A}}\log\|M_{n}(\omega,z;\xi+iy)\|d\xi dy\leq(\pi r^{2})^{-d}C(z)n\mathrm{mes}(\mathcal{A})<\exp(-C_{0}n^{\sigma}/2).$ (3.17)
Combining (3.15), (3.16) and (3.17), we obtain the conclusion. ∎
Corollary 3.11.
Suppose $\omega_{0}\in\mathbb{T}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $\sigma$, $\tau$ be as in LDT. Then for all $n\geq N_{0}(p,q,z_{0},\gamma)$ and $(\omega,z,y)\in\mathbb{C}^{d}\times\partial\mathbb{D}\times\mathbb{R}^{d}$ such that
$|y|<\frac{1}{n},\quad|\omega-\omega_{0}|,\quad|z-z_{0}|<\exp(-(\log n)^{8/\sigma}),$
we have
$\sup\limits_{x\in\mathbb{T}^{d}}\log\|M_{n}(\omega,z;x+iy)\|\leq nL_{n}(\omega_{0},z_{0})+Cn^{1-\tau},$
where $C=C(p,q,z_{0},\gamma)$.
Proof.
Take $\omega,z,y$ satisfying the assumptions. Let $A=2/\sigma$. According to Lemma 3.10 and Lemma 3.5, we obtain
$\log\|M_{n}(\omega,z;x+iy)\|\leq nL_{n}(\omega_{0},z_{0};y)+Cn^{1-\tau}$
for all $x$ outside a set $B_{y}\in\mathbb{T}^{d}$ with $\mathrm{mes}(B_{y})<\exp(-(\log n)^{2})$.
Similar to the proof of Lemma 3.10, we can get the conclusion. ∎
Corollary 3.12.
Assume $\omega_{0}\in\mathbb{T}(p,q)$, $z_{0}\in\partial\mathbb{D}$, and $L(\omega_{0},z_{0})>\gamma>0$. Let $\sigma,\tau$ be as in LDT. For all $n>N_{0}(p,q,z_{0},\gamma)$ and $(\omega_{i},z_{i},x_{i})\in\mathbb{C}^{d}\times\partial\mathbb{D}\times\mathbb{R}^{d}$ such that
$|\mathrm{Im}x_{i}|<\frac{1}{n},\,|\omega_{i}-\omega_{0}|,\,|z_{i}-z_{0}|<\exp(-(\log n)^{8/\sigma}),\,i=1,2,$
we have
$\|M_{n}(\omega_{1},z_{1};x_{1})-M_{n}(\omega_{2},z_{2};x_{2})\|\leq(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|)\exp(nL(\omega_{0},z_{0})+Cn^{1-\tau}),$
where $C=C(p,q,z_{0},\gamma)$. Particularly,
$\big{|}\log\|M_{n}(\omega_{1},z_{1};x_{1})\|-\log\|M_{n}(\omega_{2},z_{2};x_{2})\|\big{|}\leq(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|)\exp(nL(\omega_{0},z_{0})+Cn^{1-\tau}),$ (3.18)
provided the right-hand side of (3.18) is less than $\frac{1}{2}$.
Proof.
Similar to the proof of Lemma 3.4, based on the Corollary 3.11, the conclusion follows. ∎
3.3 LDT Estimate for the Entries of the Monodromy
Define the unitary matrices
$\Theta_{n}=\left[\begin{array}[]{cc}\overline{\alpha}_{n}&\rho_{n}\\ \rho_{n}&-\alpha_{n}\end{array}\right].$
Then it is valid to factorize the matrix $\mathcal{C}$ as follows,
$\mathcal{C}=\mathcal{L}_{+}\mathcal{M}_{+},$
where
$\mathcal{L}_{+}=\left[\begin{matrix}\Theta_{0}&~{}&~{}\\ ~{}&\Theta_{2}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right],\quad\mathcal{M}_{+}=\left[\begin{matrix}\mathbf{1}&~{}&~{}\\ ~{}&\Theta_{1}&~{}\\ ~{}&~{}&\ddots\end{matrix}\right].$
Similarly, the extended CMV matrix can be written as
$\mathcal{E}=\mathcal{L}\mathcal{M},$
where
$\mathcal{L}=\bigoplus_{j\in\mathbb{Z}}\Theta_{2j},\quad\mathcal{M}=\bigoplus_{j\in\mathbb{Z}}\Theta_{2j+1}.$
We let $\mathcal{E}_{[a,b]}$ denote the restriction of an extended CMV matrix to the finite interval $[a,b]$, defined by
$\mathcal{E}_{[a,b]}=P_{[a,b]}\mathcal{E}(P_{[a,b]})^{*},$
where $P_{[a,b]}$ is the projection $\ell^{2}(\mathbb{Z})\rightarrow\ell^{2}([a,b])$. $\mathcal{L}_{[a,b]}$ and $\mathcal{M}_{[a,b]}$ are defined similarly.
However, the matrix $\mathcal{E}_{[a,b]}$ will no longer be unitary due to the fact that $|\alpha_{a}|<1$ and $|\alpha_{b}|<1$. Thus, we need to modify the boundary conditions briefly. With $\beta,\eta\in\partial\mathbb{D}$, define the sequence of Verblunsky coefficients
$\hat{\alpha}_{n}=\begin{cases}\beta,\quad&n=a;\\ \eta,&n=b;\\ \alpha_{n},&n\notin\{a,b\}.\end{cases}$
Without loss of generality, we restrict the extended CMV matrix to the interval $[0,n-1]$. We denote the characteristic determinant of matrix [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}$ as $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$. According to the results in [26, Theorem 2], the relation between this characteristic determinant and the $n$-step transfer matrix is
$M_{n}(\omega,z;x)=(\sqrt{z})^{-n}\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}\left(\begin{matrix}z\varphi^{\beta,\eta}_{[1,n-1]}&\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\\ z\big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\big{)}^{*}&(\varphi^{\beta,\eta}_{[1,n-1]})^{*}\end{matrix}\right).$ (3.19)
In order to obtain the large deviation theorem for the entries of the monodromy matrix, we need to estimate the above characteristic determinant.
Remark 3.13.
According to the above matrix, we have the following two conclusions
$\sup\limits_{x\in\mathbb{T}^{d}}\log|\varphi^{\beta,\eta}_{[0,n-1]}(x+iy)|\leq nL_{n}(\omega_{0},z_{0})+Cn^{1-\tau}$
and
$\big{|}\log|\varphi^{\beta,\eta}_{[0,n-1]}(\omega_{1},z_{1};x_{1})|-\log|\varphi^{\beta,\eta}_{[0,n-1]}(\omega_{2},z_{2};x_{2})|\big{|}$
$\leq(|\omega_{1}-\omega_{2}|+|z_{1}-z_{2}|+|x_{1}-x_{2}|)\frac{\exp(nL(\omega_{0},z_{0})+Cn^{1-\tau})}{\mathop{\max}\limits_{i}|\varphi^{\beta,\eta}_{[1,n-1]}(\omega_{i},z_{i};x_{i})|}$ (3.20)
provided the right-hand side of (3.20) is less than $\frac{1}{2}$, which can be derived from Corollary 3.11 and Corollary 3.12.
Lemma 3.14.
[14, Lemma 11.4] Suppose $V$ is a nonconstant real-analytic function on $Q_{0}=[-2,2]^{d}$ with $\sup\limits_{Q_{0}}|V|\leq 1$. Then there exists $\varepsilon=\varepsilon(V,d)>0$ and $C=C(V,d)$ so that
$\mathrm{mes}\{(x_{1},\ldots,x_{d})\in[-1,1]^{d}:|V(x_{1},\ldots,x_{d})-E|<t\}\leq Ct^{\varepsilon}$
for all $-1\leq E\leq 1$ and $10<t<$ 1.
Lemma 3.15.
Fix some $z$ and let $\varphi_{[0,n-1]}^{\beta,\eta}=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$. Then
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\varphi_{[0,n-1]}^{\beta,\eta}|\leq\exp(-n^{d+2})\}\leq\exp(-n)$ (3.21)
for large $n$.
Proof.
Suppose (3.21) fails and set $u(x)=\log\big{|}\varphi_{[0,n-1]}^{\beta,\eta}\big{|}$. According to Lemma 7.2 with $M\leq C(\alpha_{n},z)n$, $L=n^{d+2}$ and $\delta=\exp(-n)$, then we have
$\sup\limits_{\mathbb{T}^{d}}u\leq C_{\rho,d}C(\alpha_{n},z)n-\frac{n^{d+2}}{C_{\rho,d}\log^{d}(C_{d}/\delta)}\leq-C_{1}n^{2}$
for some small constant $C_{1}$. Thus,
$\sup\limits_{\mathbb{T}^{d}}\big{|}\varphi_{[0,n-1]}^{\beta,\eta}\big{|}\leq\exp(-C_{1}n^{2}).$ (3.22)
On the other hand, we can write $\varphi_{[0,n-1]}^{\beta,\eta}$ as $\det(D_{[0,n-1]}+B_{[0,n-1]})$, where
$D_{[0,n-1]}=\mathrm{diag}\{z+\overline{\alpha}_{0}\alpha_{-1},z+\overline{\alpha}_{1}\alpha_{0},\ldots,z+\overline{\alpha}_{n-1}\alpha_{n-2}\}.$
Due to $|\alpha_{j}|,\rho_{j}<1$, we know that $\|B_{[0,n-1]}\|<3$.
Therefore,
$\Big{|}\frac{1}{n}\log\big{|}\varphi_{[0,n-1]}^{\beta,\eta}\big{|}\Big{|}$ $=\Big{|}\frac{1}{n}\log\big{|}\det(D_{[0,n-1]}+B_{[0,n-1]})\big{|}\Big{|}$
$\leq\Big{|}\frac{1}{n}\log\big{|}\det D_{[0,n-1]}\big{|}\Big{|}+\Big{|}\frac{1}{n}\log\big{|}\det(I+D^{-1}_{[0,n-1]}B_{[0,n-1]})\big{|}\Big{|}$
$=\Big{|}\frac{1}{n}\sum_{j=0}^{n-1}\log|z+\overline{\alpha}_{j}\alpha_{j-1}|\Big{|}+\Big{|}\frac{1}{n}\log\big{|}\det(I+D^{-1}_{[0,n-1]}B_{[0,n-1]})\big{|}\Big{|}.$
Let $\hat{\alpha}:=\sup\limits_{j\in\mathbb{Z},\,x\in\mathbb{T}^{d}}|\overline{\alpha}_{j}\alpha_{j-1}|$. Now we consider the following two cases.
Case 1. Suppose $\hat{\alpha}\geq\frac{1}{2}$. According to Lemma 3.14, we have that
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\overline{\alpha}_{j}\alpha_{j-1}-1|<e^{-\rho}\}<Ce^{-\varepsilon\rho}$
for some $\rho>0$.
Since
$\|D^{-1}_{[0,n-1]}\|\leq N_{0}^{-M}(z+\overline{\alpha}_{j}\alpha_{j-1})|^{-1},$
we get
$\mathrm{mes}\Big{\{}x\in\mathbb{T}^{d}:\|D^{-1}_{[0,n-1]}\|>\frac{2}{3}\Big{\}}$ $\leq n\mathrm{mes}\Big{\{}x\in\mathbb{T}^{d}:|N_{0}^{-M}(z+\overline{\alpha}(Tx)\alpha(x))|<\frac{3}{2}N_{0}^{-M}\Big{\}}$
$=n\mathrm{mes}\Big{\{}x\in\mathbb{T}^{d}:|N_{0}^{-M}z^{-1}\overline{\alpha}(Tx)\alpha(x)-N_{0}^{-M}|<\frac{3}{2}N_{0}^{-M}\Big{\}}$
$\leq CnN_{0}^{-\varepsilon M}.$
Thus,
$\mathrm{mes}\{x\in\mathbb{T}^{d}:\|D^{-1}_{[0,n-1]}B_{[0,n-1]}\|\geq 2\}\leq CnN_{0}^{-\varepsilon M}.$
Then we have
$\Big{|}\frac{1}{n}\log\big{|}\varphi_{[0,n-1]}^{\beta,\eta}\big{|}\Big{|}$ $\leq\Big{|}\frac{1}{n}\log\big{|}\det D_{[0,n-1]}\big{|}\Big{|}+\Big{|}\frac{1}{n}\log\big{|}\det(I+D^{-1}_{[0,n-1]}B_{[0,n-1]})\big{|}\Big{|}$
$\leq\log(1+\hat{\alpha})+\log 3=\log(3(1+\hat{\alpha}))=:C(\hat{\alpha}).$
That is, $e^{-C(\hat{\alpha})n}\leq|\varphi_{[0,n-1]}^{\beta,\eta}|\leq e^{C(\hat{\alpha})n}$, which contradicts (3.22).
Case 2. Suppose $\hat{\alpha}<\frac{1}{2}$. According to the definition of $D_{[0,n-1]}$,
$\Big{|}\frac{1}{n}\log|\det D_{[0,n-1]}|\Big{|}=\Big{|}\frac{1}{n}\sum_{j=0}^{n-1}\log|z+\overline{\alpha}_{j}\alpha_{j-1}|\Big{|}.$
If $|z+\overline{\alpha}_{j}\alpha_{j-1}|<1$, then
$|z+\overline{\alpha}_{j}\alpha_{j-1}|\geq|z|-|\overline{\alpha}_{j}\alpha_{j-1}|\geq|z|-\hat{\alpha}\geq\frac{1}{2}.$
If $|z+\overline{\alpha}_{j}\alpha_{j-1}|>1$, then
$|z+\overline{\alpha}_{j}\alpha_{j-1}|\leq|z|+|\overline{\alpha}_{j}\alpha_{j-1}|\leq 2.$
Thus, $\Big{|}\frac{1}{n}\log|\det D_{[0,n-1]}|\Big{|}<\log 2<2$ and
$\Big{|}\frac{1}{n}\log\big{|}\varphi_{[0,n-1]}^{\beta,\eta}\big{|}\Big{|}$ $\leq\Big{|}\frac{1}{n}\log\big{|}\det D_{[0,n-1]}\big{|}\Big{|}+\Big{|}\frac{1}{n}\log\big{|}\det(I+D^{-1}_{[0,n-1]}B_{[0,n-1]})\big{|}\Big{|}$
$\leq 2+\log 7<4.$
This implies that $e^{-4n}\leq|\varphi_{[0,n-1]}^{\beta,\eta}|\leq e^{4n}$, which contradicts (3.22).
In conclusion, $\mathrm{mes}\{x\in\mathbb{T}^{d}:|\varphi_{[0,n-1]}^{\beta,\eta}|\leq\exp(-n^{d+2})\}\leq\exp(-n)$. ∎
Lemma 3.16.
Let $N_{0}$ be as in Lemma 3.15 and $10<\theta<$ 1. Fix some $z$ and let $\varphi_{[0,n-1]}^{\beta,\eta}=\det\big{(}z-\mathcal{E}^{\beta,\eta}_{[0,n-1]}\big{)}$. Then for any $N_{0}\leq n<N^{\frac{1-\theta}{2}}$,
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\varphi_{[0,n-1]}^{\beta,\eta}|\leq\exp(-N^{1-\theta})\}\leq\exp(-N^{\frac{1-\theta}{d}}n^{-\frac{2}{d}}),$ (3.23)
where $N$ is large.
Proof.
Suppose (3.23) fails for some $n$ and $N$. Applying Lemma 7.2 with $M=n$, $\delta=\exp(-N^{\frac{1-\theta}{d}}n^{-\frac{2}{d}})$ and $L=N^{1-\theta}$, we obtain
$\sup\limits_{\mathbb{T}^{d}}u\leq Cn-\frac{N^{1-\theta}}{C\log^{d}(C/\theta)}\leq-C_{1}n^{2}$
for some constant $C_{1}$. In other words,
$\sup\limits_{\mathbb{T}^{d}}|\varphi_{[0,n-1]}^{\beta,\eta}|\leq\exp(-C_{1}n^{2}).$
Then this leads to a contradiction in the same way as the derivation of the contradiction in the proof of Lemma 3.15. Similar to the proof of Lemma 3.15, the conclusion follows. ∎
Now we introduce the properties of matrices in $SL(2,\mathbb{C})$. For any $A\in SL(2,\mathbb{C})$, there are unit vectors $\underline{u}^{+}_{A}$, $\underline{u}^{-}_{A}$, $\underline{v}^{+}_{A}$, $\underline{v}^{-}_{A}$ such that $A\underline{u}^{+}_{A}=\|A\|\underline{v}^{+}_{A}$, $A\underline{u}^{-}_{A}=\|A\|^{-1}\underline{v}^{-}_{A}$. Moreover, $\underline{u}^{+}_{A}\bot\underline{u}^{-}_{A}$ and $\underline{v}^{+}_{A}\bot\underline{v}^{-}_{A}$. In what follows, we denote that $a\wedge b:=\min\{a,b\}$ for positive numbers $a,b$ and $\varphi\wedge\psi$ stands for the exterior product of vectors $\varphi,\psi$.
Lemma 3.17.
From the Lemma 3.2, we have
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x)\|-nL_{n}(\omega,z)|>n^{1-\tau}\}<\exp(-C_{0}n^{\sigma}).$ (3.24)
Then
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\varphi_{[0,n-1]}^{\beta,\eta}(x)|+|\varphi_{[0,n-1]}^{\beta,\eta}(T^{j_{1}}x)|+|\varphi_{[0,n-1]}^{\beta,\eta}(T^{j_{2}}x)|\leq\exp(nL_{n}(\omega,z)-100n^{1-\sigma})\}$
$\leq\exp(-n^{\frac{1-\sigma}{2d}}\wedge\sigma)$ (3.25)
for any $l_{0}=l_{0}(T,z,\sigma)\leq j_{1}\leq j_{1}+l_{0}\leq j_{2}\leq n^{\frac{1-\sigma}{8}}$ and $n>N(T,C_{0},z,\sigma)$. Moreover, to obtain (3.25) for some $n$ only requires (3.24) with the same $n$.
Proof.
From (3.19), the transfer matrix $M_{n}(\omega,z;x)\in SU(1,1)$, then we can verify that
$T_{n}(\omega,z;x)=Q^{*}M_{n}(\omega,z;x)Q=(\sqrt{z})^{-n}\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}\left(\begin{matrix}t_{1}&t_{2}\\ t_{3}&t_{4}\end{matrix}\right)\in SL(2,\mathbb{R}),$ (3.26)
where
$t_{1}=\frac{1}{2}z\varphi^{\beta,\eta}_{[1,n-1]}+\frac{1}{2}z\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*}+\frac{1}{2}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}+\frac{1}{2}(\varphi^{\beta,\eta}_{[1,n-1]})^{*},$
$t_{2}=-\frac{i}{2}z\varphi^{\beta,\eta}_{[1,n-1]}-\frac{i}{2}z\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*}+\frac{i}{2}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}+\frac{i}{2}(\varphi^{\beta,\eta}_{[1,n-1]})^{*},$
$t_{3}=\frac{i}{2}z\varphi^{\beta,\eta}_{[1,n-1]}-\frac{i}{2}z\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*}+\frac{i}{2}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}-\frac{i}{2}(\varphi^{\beta,\eta}_{[1,n-1]})^{*},$
$t_{4}=\frac{1}{2}z\varphi^{\beta,\eta}_{[1,n-1]}-\frac{1}{2}z\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*}-\frac{1}{2}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}+\frac{1}{2}(\varphi^{\beta,\eta}_{[1,n-1]})^{*}$
and “*” stands for the Szegő dual of the corresponding polynomial.
According to [26, Section 3], we know that
$\varphi^{\beta,\eta}_{[0,n-1]}=(z+\overline{\alpha}_{0}\alpha_{-1})\varphi^{\beta,\eta}_{[1,n-1]}-\rho_{0}\alpha_{-1}\det\mathcal{P}_{n-1},$ (3.27)
where
$\mathcal{P}_{n-1}=\left[\begin{array}[]{ccccccc}-\overline{\alpha}_{1}\rho_{0}&-\rho_{1}\rho_{0}&&&&\\ -\overline{\alpha}_{2}\rho_{1}&z+\overline{\alpha}_{2}\alpha_{1}&-\overline{\alpha}_{3}\rho_{2}&-\rho_{3}\rho_{2}&&\\ -\rho_{2}\rho_{1}&\rho_{2}\alpha_{1}&z+\overline{\alpha}_{3}\alpha_{2}&\rho_{3}\alpha_{2}&&\\ &&-\overline{\alpha}_{4}\rho_{3}&z+\overline{\alpha}_{4}\alpha_{3}&&\\ &&&&\ddots&\\ &&&&&z+\overline{\eta}\alpha_{n-2}\end{array}\right].$
Then for unit vectors $\underline{e}_{1}=(1,0)^{T}$ and $\underline{e}_{2}=(0,1)^{T}$, we can derive
$T_{n}(\omega,z;x)\underline{e}_{1}\wedge\underline{e}_{2}=(\sqrt{z})^{-n}\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}t_{1}.$
Due to the fact that $(z\varphi^{\beta,\eta}_{[1,n-1]})^{*}=(\varphi^{\beta,\eta}_{[1,n-1]})^{*}$ and
$\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*}=z\Big{(}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\eta}_{[0,n-1]}}{\alpha_{-1}}\Big{)}^{*},$
we have
$|T_{n}(\omega,z;x)\underline{e}_{1}\wedge\underline{e}_{2}|$ $\leq\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}\Big{(}|\varphi^{\beta,\eta}_{[1,n-1]}|+\Big{|}\frac{z\varphi^{\beta,\eta}_{[1,n-1]}-\varphi^{\beta,\gamma}_{[0,n-1]}}{\alpha_{-1}}\Big{|}\Big{)}$
$=\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}\Big{(}|\varphi^{\beta,\eta}_{[1,n-1]}|+|\rho_{0}\det\mathcal{P}_{n-1}-\overline{\alpha}_{0}\varphi^{\beta,\eta}_{[1,n-1]}|\Big{)}$
$\leq\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}(2|\varphi^{\beta,\eta}_{[1,n-1]}|+|\det\mathcal{P}_{n-1}|).$
Therefore, from (3.27) we have
$|\varphi^{\beta,\eta}_{[0,n-1]}|\leq 2|\varphi^{\beta,\eta}_{[1,n-1]}|+|\det\mathcal{P}_{n-1}|\leq\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}(2|\varphi^{\beta,\eta}_{[1,n-1]}|+|\det\mathcal{P}_{n-1}|).$
Suppose
$|T_{n}(\omega,z;x)\underline{e}_{1}\wedge\underline{e}_{2}|\leq\Big{(}\prod_{j=0}^{n-1}\frac{1}{\rho_{j}}\Big{)}(2|\varphi^{\beta,\eta}_{[1,n-1]}|+|\det\mathcal{P}_{n-1}|)<\exp(nL_{n}(z)-100n^{1-\sigma}).$
For convenience, we write $T_{n}(\omega,z;x)$ as $T_{n}$. Since
$T_{n}\underline{u}^{+}_{n}\wedge\underline{e}_{2}(\underline{u}^{+}_{n}\cdot\underline{e}_{1})+T_{n}\underline{u}^{-}_{n}\wedge\underline{e}_{2}(\underline{u}^{-}_{n}\cdot\underline{e}_{1})=\|T_{n}\|\underline{v}^{+}_{n}\wedge\underline{e}_{2}(\underline{u}^{+}_{n}\cdot\underline{e}_{1})+\|T_{n}\|^{-1}\underline{v}^{-}_{n}\wedge\underline{e}_{2}(\underline{u}^{-}_{n}\cdot\underline{e}_{1}),$
one can obtain that
$\|T_{n}\||\underline{u}^{+}_{n}\cdot\underline{e}_{1}||\underline{v}^{+}_{n}\wedge\underline{e}_{2}|-\|T_{n}\|^{-1}|\underline{u}^{-}_{n}\cdot\underline{e}_{1}||\underline{v}^{-}_{n}\wedge\underline{e}_{2}|\leq\exp(nL_{n}(z)-100n^{1-\sigma}).$
Then the proof of (3.25) is the same as the proof of the Lemma 2.9 in [13], we just need to replace the use of $f_{N}$ there by our $\varphi^{\beta,\eta}_{[0,n-1]}$ and replace $N$ by $n$. ∎
Lemma 3.18.
There exists some constant $\kappa>0$ (depending on $\sigma$ in LDT) such that
$\int_{\mathbb{T}^{d}}\frac{1}{n}\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|dx>L_{n}(z)-n^{-\kappa}$ (3.28)
for $n\geq N(T,z,\sigma)$.
Proof.
Replace $f_{N}(x)$ in the proof of Lemma 2.10 in [13] by $\varphi^{\beta,\eta}_{[0,n-1]}(x)$, with the aid of the Lemma 3.16 and Lemma 3.17, one can get the statement. ∎
Lemma 3.19.
From the Lemma 3.10, we have the upper bound
$\sup\limits_{\mathbb{T}^{d}}\log\|M_{n}(\omega,z;x)\|\leq nL_{n}(\omega,z)+Cn^{1-\tau}.$ (3.29)
Fix a large positive integer $n\geq N_{0}$, then for some small constant $\nu$ depending on $\sigma$ and $\tau$,
$\mathrm{mes}\Big{\{}x\in\mathbb{T}^{d}:\frac{1}{n}\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|<L_{n}(z)-n^{-\nu}\Big{\}}\leq\exp(-Cn^{\nu}).$ (3.30)
Proof.
We just consider the case that $d=2$. The general case is similar.
[…]
$\langle u\rangle>L_{n}-n^{-\kappa},\quad\sup\limits_{\mathbb{T}^{2}}u<L_{n}+Cn^{-\tau}.$
[…]
$\|g-\langle g\rangle\|_{1}\leq Cn^{-\nu}$ (3.31)
for some constant $C$ and $\nu$ (smaller than $\tau$ and $\kappa$).
Since $\sup\limits_{\mathbb{T}}g\geq\langle g\rangle>\frac{\gamma}{2}>0$ for large $n$ and $\sup\limits_{T^{2}_{h}}u\leq C$, one can conclude from the proof of Lemma 7.2 that there is the Riesz representation
$g(z)=\int_{\mathbb{T}_{h/2}}\log|z-\zeta|d\mu(\zeta)+h(z),$
where
$\mu(\mathbb{T}_{h/2})+\|h\|_{L^{\infty}(\mathbb{T}_{h/2})}\leq C_{h}(m+M),$
provided that $\sup\limits_{y\in\mathbb{T}}u(z,y)>-m$.
According to the definition of the norm on BMO($\mathbb{T}$), one can derive that
$\|g\|_{BMO(\mathbb{T})}\leq C\|g-\langle g\rangle\|^{1/2}_{1}\leq Cn^{-\frac{\nu}{2}}.$ (3.32)
[…]
$\mathrm{mes}(\mathbb{T}\backslash\Omega)\leq\exp(-n^{\nu/4}).$ (3.33)
Fix some $x\in\Omega$. Then
[…]
Since $\sup\limits_{y\in\mathbb{T}}u(x,y)\geq 0$ and $\sup\limits_{z\in\mathbb{T}_{h}}u(x,z)\leq C$, combining (3.34) and the definition of the norm on BMO($\mathbb{T}$), we have
$\|u(x,\cdot)\|_{BMO(\mathbb{T})}\leq Cn^{-\nu/8}$
for any $x\in\Omega$.
Therefore, from the John-Nirenberg inequality,
$\mathrm{mes}\{y\in\mathbb{T}|u(x,y)<L_{n}-Cn^{-\nu/16}\}\leq\exp(-Cn^{\nu/16})$ (3.35)
for large $n$. Then the statement follows from (3.33) and (3.35) via Fubini. ∎
Remark 3.20.
From the above lemma, we get the LDT for the determinant $\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)$ as follows
$\mathrm{mes}\{x\in\mathbb{T}^{d}:|\log|\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)|-nL_{n}(\omega,z)|>n^{1-\tau}\}<\exp(-Cn^{\nu}).$ (3.36)
In what follows, we denote $\mathcal{D}(z_{0},r)=\{z\in\mathbb{C}:|z-z_{0}|<r\}$.
Lemma 3.21.
Suppose that $x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$ and $L(\omega,z)>\gamma>0$. Let $H\gg 1$ and $\tau$, $\nu$ as in LDT. There exists $C_{0}=C_{0}(p,q,z,\gamma)$, such that for any $n\geq N(p,q,z,\gamma)$, if
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|\leq nL_{n}(\omega,z)-C_{0}Hn^{1-\tau},$
then there exists $x^{\prime}\in\mathbb{C}^{d}$, $|x^{\prime}-x|\lesssim\exp(-(H+n^{\nu}/d))$ such that $\varphi^{\beta,\eta}_{[0,n-1]}(x^{\prime})=0$. Furthermore,
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-z)^{-1}\|\geq C\exp(H+n^{\nu}/d)$
for some $C$.
Proof.
From Remark 3.20, there exists $x_{0}$ outside the set in (3.36) with $|x-x_{0}|\lesssim\exp(-n^{\nu}/d)$ such that
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|>nL_{n}(\omega,z)-n^{1-\tau}=:m.$
Let
$\phi(\zeta)=\varphi^{\beta,\eta}_{[0,n-1]}\Big{(}x_{0}+\frac{C\exp(-n^{\nu}/d)\zeta}{|x-x_{0}|}(x-x_{0})\Big{)}$
and $\zeta_{0}\in\mathcal{D}(0,\frac{1}{7})$ such that $\phi(\zeta_{0})=\varphi^{\beta,\eta}_{[0,n-1]}(x)$.
From the assumption, one has
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|\leq nL_{n}(\omega,z)-C_{0}Hn^{1-\tau}<nL_{n}(\omega,z)=:M.$
According to the Cartan’s estimate, there exists a set $\mathcal{B}\in\mathrm{Car}_{1}(H,CHn^{1-\tau})$ such that
$\log|\phi(\zeta)|>M-CH(M-m)=CHn^{1-\tau}$
for all $\zeta\in\mathcal{D}(0,\frac{1}{6})\backslash\mathcal{B}$.
On the other hand,
$\log|\phi(\zeta_{0})|=\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|\leq nL_{n}(\omega,z)-C_{0}Hn^{1-\tau}.$
Thus, there exists $\zeta_{j}\in\mathcal{B}$ such that $\zeta_{0}\in\mathcal{D}(\zeta_{j},r_{j})\subset\mathcal{B}$, where $r_{j}<\exp(-H)$. Then Lemma 7.3 implies that there exists $\zeta^{\prime}\in\mathcal{D}(\zeta_{j},r_{j})$ such that $\phi(\zeta^{\prime})=0$.
Taking $x^{\prime}=x_{0}+\frac{C\exp(-n^{\nu}/d)\zeta^{\prime}}{|x-x_{0}|}(x-x_{0})$, then we can obtain $\varphi^{\beta,\eta}_{[0,n-1]}(x^{\prime})=0$.
Due to the structure of the matrix [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}$, there exists some constant $C$ such that
[$\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|\leq C|x-x^{\prime}|.$
In addition, if
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|<1,$
then we have
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})\|$
=([$=\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}\|\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)\|$
$<1$
and
$\|E-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)\|<1,$
where $E$ denotes the identity matrix.
Thus, $E-[E-(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x^{\prime})-E)^{-1}(\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E)]$ is invertible and this implies that [$\mathcal{E}^{\beta,\eta}_{[0,n-1]}(\omega,z;x)-E$ is invertible. Then the second statement follows. ∎
As a consequence, we have the spectral form of LDT.
Lemma 3.22.
Assume $x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$ and $L(\omega,z)>\gamma>0$. Let $\tau$, $\nu$ as in LDT. If $n\geq N(p,q,z,\gamma)$ and
([$\|(\mathcal{E}^{\beta,\eta}_{[0,n-1]}-z)^{-1}\|\leq C\exp(n^{\nu/2}),$
then
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|>nL_{n}(\omega,z)-n^{1-\tau/2}.$
Proof.
Let $n^{\nu/2}=H+\frac{n^{\nu/2}}{d}$, then $H=\frac{d-1}{d}n^{\nu/2}$. Applying the above lemma, one can obtain
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(x)|$ $>nL_{n}(\omega,z)-C_{0}Hn^{1-\tau}=nL_{n}(\omega,z)-C_{0}\frac{d-1}{d}n^{\nu/2}n^{1-\tau}$
$>nL_{n}(\omega,z)-C_{0}\frac{d-1}{d}n^{\nu/2}n^{\frac{1}{2}-\frac{\tau}{2}}>nL_{n}(\omega,z)-C_{0}\frac{d-1}{d}n^{1-\frac{\tau}{2}}$
$>nL_{n}(\omega,z)-n^{1-\frac{\tau}{2}}.$
3.4 Poisson Formula for CMV Matrix
For $z\in\mathbb{C}$, $\beta,\eta\in\partial\mathbb{D}$, we can define the polynomials
$\varphi^{\beta,\eta}_{[a,b]}(z):=\det(z-\mathcal{E}^{\beta,\eta}_{[a,b]}),\quad\phi^{\beta,\eta}_{[a,b]}(z):=(\rho_{a}\cdots\rho_{b})^{-1}\varphi^{\beta,\eta}_{[a,b]}(z).$
Note that when $a>b$, $\phi^{\beta,\eta}_{[a,b]}(z)=1$.
Since the equation $\mathcal{E}u=zu$ is equivalent to $(z\mathcal{L}^{*}-\mathcal{M})u=0$. Then the associated finite-volume Green’s functions are as follows
$G^{\beta,\eta}_{[a,b]}(z)=\big{(}z(\mathcal{L}^{\beta,\eta}_{[a,b]})^{*}-\mathcal{M}^{\beta,\eta}_{[a,b]}\big{)}^{-1},$
$G^{\beta,\eta}_{[a,b]}(j,k;z)=\langle\delta_{j},G^{\beta,\eta}_{[a,b]}(z)\delta_{k}\rangle,\quad j,k\in[a,b].$
According to [17, Proposition 3.8] and [30, Section B.1], for $\beta,\eta\in\partial\mathbb{D}$, the Green’s function has the following expression
$|G^{\beta,\eta}_{[a,b]}(j,k;z)|=\frac{1}{\rho_{k}}\Big{|}\frac{\phi^{\beta,\cdot}_{[a,j-1]}(z)\phi^{\cdot,\eta}_{[k+1,b]}(z)}{\phi^{\beta,\eta}_{[a,b]}(z)}\Big{|},\quad a\leq j\leq k\leq b,$
where $\cdot$” stands for the unchanged Verblunsky coefficient.
From [17, Lemma 3.9], if $u$ satisfies $\mathcal{E}u=zu$, Poisson formula reads
$u(m)=$ $G^{\beta,\eta}_{[a,b]}(a,m;z)\begin{cases}(z\overline{\beta}-\alpha_{a})u(a)-\rho_{a}u(a+1),\quad&a\text{ even,}\\ (z\alpha_{a}-\beta)u(a)+z\rho_{a}u(a+1),&a\text{ odd,}\end{cases}$
$+G^{\beta,\eta}_{[a,b]}(m,b;z)\begin{cases}(z\overline{\eta}-\alpha_{b})u(b)-\rho_{b}u(b-1),\quad&b\text{ even,}\\ (z\alpha_{b}-\eta)u(b)+z\rho_{b-1}u(b-1),&b\text{ odd}\end{cases}$
for $a<m<b$.
Lemma 3.23.
Let $x,\omega\in\mathbb{T}^{d}$, $z\in\partial\mathbb{D}$ and $[a,b]\subset\mathbb{Z}$. If for any $m\in[a+2,b-2]$, there exists an interval $I_{m}=[a_{m},b_{m}]\subset[a+2,b-2]$ containing $m$ such that
$|G^{\beta,\eta}_{I_{m}}(a_{m},m;z)|\begin{cases}|z\overline{\beta}-\alpha_{a}|+\rho_{a},&a\text{ even,}\\ |z\alpha_{a}-\beta|+\rho_{a},&a\text{ odd,}\end{cases}+|G^{\beta,\eta}_{I_{m}}(m,b_{m};z)|\begin{cases}|z\overline{\eta}-\alpha_{b}|+\rho_{b},&b\text{ even}\\ |z\alpha_{b}-\eta|+\rho_{b-1},&b\text{ odd}\end{cases}<1,$
then $z\notin\sigma(\mathcal{E}^{\beta,\eta}_{[a,b]})$.
Proof.
Assume to the contrary that $z\in\sigma(\mathcal{E}^{\beta,\eta}_{[a,b]})$ and let $u$ be a corresponding eigenvector, i.e., [$\mathcal{E}^{\beta,\eta}_{[a,b]}u=zu$. Suppose that $|u(m)|=\max\limits_{n\in(a,b)}|u(n)|$. If $m\in[a+2,b-2]$, then there is a vector $\tilde{u}$ satisfying [$\mathcal{E}^{\beta,\eta}_{[a,b]}\tilde{u}=z\tilde{u}$ and $\tilde{u}|_{[a+2,b-2]}=u|_{[a+2,b-2]}$. Thus, from the hypothesis and the Poisson formula, one can get that
$|u(m)|\leq\max\{|u(a_{m})|,|u(a_{m}+1)|,|u(b_{m})|,|u(b_{m}-1)|\},$
which is a contradiction. ∎
Remark 3.24.
In the above covering lemma, the choice of the interval $[a+2,b-2]$ depends on the structure of the matrix [$\mathcal{E}^{\beta,\eta}_{[a,b]}$. Both this matrix and Poisson formula depend on the parity of $a$ and $b$. In addition, the purpose of the boundary condition $\beta$ and $\eta$ is to guarantee that the restriction of matrix $\mathcal{E}$ to $[a,b]$ is unitary. In fact, no matter how we choose $a$ and $b$ (the parity of $a$ and $b$), there are at most the first two rows and the last two rows relevant to the parameters $\beta$ and $\eta$. But under the condition of parameters $\beta$ and $\eta$, Poisson formula is no longer applicable. Thus, by considering the interval $[a+2,b-2]$ and applying the Poisson formula, we can prove the above covering lemma.
Lemma 3.25.
Suppose that $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $K\in\mathbb{R}$ and $\tau$ be as in LDT. There exists $C_{0}=C_{0}(p,q,z_{0},\gamma)$ such that if $n\geq N(p,q,z_{0},\gamma)$ and
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(\omega_{0},z_{0};x_{0})|>nL_{n}(\omega_{0},z_{0})-K,$ (3.37)
then for any $(\omega,z,x)\in\mathbb{T}^{d}\times\partial\mathbb{D}\times\mathbb{T}^{d}$ with $|x-x_{0}|,|\omega-\omega_{0}|,|z-z_{0}|<\exp(-(K+C_{0}n^{1-\tau}))$ we have
$|G^{\beta,\eta}_{[0,n-1]}(j,k;z)|\leq\exp(-\frac{\gamma}{2}|k-j|+K+2C_{0}n^{1-\tau})$ (3.38)
and
$\|G^{\beta,\eta}_{[0,n-1]}(z)\|\leq\exp(K+3C_{0}n^{1-\tau}).$ (3.39)
Proof.
Take $|x-x_{0}|,|\omega-\omega_{0}|,|z-z_{0}|<\exp(-(K+C_{0}n^{1-\tau}))$ with $C$ large enough.
According to Remark 3.13, we have
$\log|\varphi^{\beta,\eta}_{[0,n-1]}(\omega,z;x)|$ $\geq\log|\varphi^{\beta,\eta}_{[0,n-1]}(\omega_{0},z_{0};x_{0})|-1\geq nL_{n}(\omega_{0},z_{0})-K-1$
$\geq nL_{n}(\omega,z)-K-2\geq nL(\omega,z)-K-2.$
Thus, according to Lemma 3.10, for $10\leq j\leq k\leq n-$ 1, one can obtain
$|G^{\beta,\eta}_{[0,n-1]}(j,k;z)|$ $=\frac{1}{\rho_{k}}\Big{|}\frac{\phi^{\beta,\cdot}_{[0,j-1]}(z)\phi^{\cdot,\eta}_{[k+1,n-1]}(z)}{\phi^{\beta,\eta}_{[0,n-1]}(z)}\Big{|}$
$=\frac{1}{\rho_{k}}\Big{|}\frac{(\rho_{0}\cdots\rho_{j-1})^{-1}\varphi^{\beta,\cdot}_{[0,j-1]}(z)(\rho_{k+1}\cdots\rho_{n-1})^{-1}\varphi^{\cdot,\eta}_{[k+1,n-1]}(z)}{(\rho_{0}\cdots\rho_{n-1})^{-1}\varphi^{\beta,\eta}_{[0,n-1]}(z)}\Big{|}$
$=(\rho_{j}\cdots\rho_{k-1})\Big{|}\frac{\varphi^{\beta,\cdot}_{[0,j-1]}(z)\varphi^{\cdot,\eta}_{[k+1,n-1]}(z)}{\varphi^{\beta,\eta}_{[0,n-1]}(z)}\Big{|}$
$\leq\Big{|}\frac{\varphi^{\beta,\cdot}_{[0,j-1]}(z)\varphi^{\cdot,\eta}_{[k+1,n-1]}(z)}{\varphi^{\beta,\eta}_{[0,n-1]}(z)}\Big{|}$
$\leq\exp(-(k-j)L(\omega,z)+Cn^{1-\tau}+K+2)$
$\leq\exp(-\frac{\gamma}{2}(k-j)+K+2C_{0}n^{1-\tau}).$
Therefore, $|G^{\beta,\eta}_{[0,n-1]}(j,k;z)|\leq\exp(-\frac{\gamma}{2}|k-j|+K+2C_{0}n^{1-\tau})$ for any choice of $j,k$ and then estimate (3.39) follows. ∎
Based on the above analysis, we can obtain the covering form of LDT.
Lemma 3.26.
Suppose that $n\gg 1$, $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $\tau,\nu$ be as in LDT. Suppose that for each point $m\in[0,n-1]$, there exists an interval $I_{m}\subset[0,n-1]$ such that:
(i) $\mathrm{dist}(m,[0,n-1]\backslash I_{m})\geq|I_{m}|/100$;
(ii) $|I_{m}|\geq C(p,q,z_{0},\gamma)$;
(iii) $\log|\varphi^{\beta,\eta}_{I_{m}}(\omega_{0},z_{0};x_{0})|>|I_{m}|L_{|I_{m}|}(\omega_{0},z_{0})-|I_{m}|^{1-\tau/4}$.
Then for any $(\omega,z,x)\in\mathbb{T}^{d}\times\partial\mathbb{D}\times\mathbb{T}^{d}$ such that
$|x-x_{0}|,\,|\omega-\omega_{0}|,\,|z-z_{0}|<\exp(-2\max\limits_{m}|I_{m}|^{1-\tau/4}),$
we have
$\mathrm{dist}(z,\sigma(\mathcal{E}^{\beta,\eta}_{[0,n-1]}))\geq\exp(-2\max\limits_{m}|I_{m}|^{1-\tau/4}).$
In addition, if $\omega\in\mathbb{T}^{d}(p,q)$ and $\max\limits_{m}|I_{m}|\leq n^{\nu/2}$, then
$\log|\varphi^{\beta,\eta}_{[0,n-1]}|>nL_{n}(\omega,z)-n^{1-\tau/2}.$
Proof.
According to Lemma 3.25, from (iii) we can obtain that
$|G^{\beta,\eta}_{[0,n-1]}(m,k;z)|\leq\exp(-\frac{\gamma}{2}|m-k|+\frac{3}{2}|I_{m}|^{1-\tau/4})$
provided
$|x-x_{0}|,\,|\omega-\omega_{0}|,\,|z-z_{0}|<\exp(-\frac{3}{2}|I_{m}|^{1-\tau/4}).$
Then Lemma 3.23 implies that $z\notin\sigma(E^{\beta,\eta}_{[0,n-1]})$ for any $|z-z_{0}|<\exp(-\frac{3}{2}|I_{m}|^{1-\tau/4})$.
Thus, if $|z-z_{0}|<\exp(-\frac{3}{2}|I_{m}|^{1-\tau/4})/2$, then
$\mathrm{dist}(z,\sigma(\mathcal{E}^{\beta,\eta}_{[0,n-1]}))\geq\exp(-\frac{3}{2}|I_{m}|^{1-\tau/4})/2.$
This implies the first statement.
If $\max\limits_{m}|I_{m}|\leq n^{\nu/2}$, we have that
$\mathrm{dist}(z,\sigma(\mathcal{E}^{\beta,\eta}_{[0,n-1]}))\geq\exp(-n^{\nu/2}).$
Then the second statement can be shown according to the spectral form of LDT. ∎
3.5 Finite Scale Localization
In this section, we derive the finite scale localization independently of the elimination of resonances.
Theorem 3.27.
Assume $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $\tau,\nu$ be as in LDT and $l,n$ be integers such that $C(p,q,z_{0},\gamma)\leq l\leq n^{\nu/2}$. Denote $z_{j}^{\Lambda}(\omega,x)$, $u_{j}^{\Lambda}(\omega,x)$ be the eigenpairs of $\mathcal{E}_{\Lambda}^{\beta,\eta}$ with $u_{j}^{\Lambda}(\omega,x)$ a unit vector, where $\mathcal{E}_{\Lambda}^{\beta,\eta}$ is the restriction to a finite interval $\Lambda$ of the extended matrix $\mathcal{E}$ and $\beta,\eta\in\partial\mathbb{D}$. Suppose that there exists an interval $I=[n^{\prime},n^{\prime\prime}]\subset[0,n-1]$ such that
$\log\big{|}\varphi_{[0,l-1]}^{\beta,\eta}(\omega_{0},z_{0};x_{0}+(m-1)\omega_{0})\big{|}>lL_{l}(\omega_{0},z_{0})-l^{1-\tau/4}$ (3.40)
for any $m\in[0,n-l]\backslash I$. Then for any $(\omega,x)\in\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d}$ satisfying $|x-x_{0}|,|\omega-\omega_{0}|<\exp(-l)$ and any eigenvalue $|z_{j}^{[0,n-1]}(\omega,x)-z_{0}|<\exp(-l)$, we have
$|u_{j}^{[0,n-1]}(\omega,x;s)|<\exp(-\frac{\gamma}{12}\mathrm{dist}(s,I))$ (3.41)
provided $\mathrm{dist}(s,I)\geq l^{2/\nu}$.
Proof.
Take $x,\omega,z=z_{j}^{[0,n-1]}(\omega,x)$ satisfying the assumptions and $s\in[0,n-1]\backslash I$ such that $d=\mathrm{dist}(s,I)\geq l^{2/\nu}$. Suppose $k\in[0,n^{\prime}]$. Let $J=[s-d,s+d]\cap[0,n-1]=:[t,n^{\prime}]$. Obviously, $|J|\geq l^{2/\nu}$.
According to the covering form of LDT, we can obtain
$\log|\varphi_{J}^{\beta,\eta}(\omega,z;x)|\geq|J|L_{|J|}(\omega,z)-|J|^{1-\tau/2}.$ (3.42)
Let $u=u_{j}^{[0,n-1]}(\omega,x)$. From Poisson formula, for $s>t$,
$u(s)=$ $G^{\beta,\eta}_{J}(s,t;z)\begin{cases}(z\overline{\beta}-\alpha_{t})u(t)-\rho_{t}u(t+1),\quad&t\text{ even,}\\ (z\alpha_{t}-\beta)u(t)+z\rho_{t}u(t+1),&t\text{ odd,}\end{cases}$
$+G^{\beta,\eta}_{J}(s,n^{\prime};z)\begin{cases}(z\overline{\eta}-\alpha_{n^{\prime}})u(n^{\prime})-\rho_{n^{\prime}}u(n^{\prime}-1),\quad&n^{\prime}\text{ even,}\\ (z\alpha_{n^{\prime}}-\eta)u(n^{\prime})+z\rho_{n^{\prime}-1}u(n^{\prime}-1),&n^{\prime}\text{ odd}.\end{cases}$
From Lemma 3.25 we have
$|G^{\beta,\eta}_{J}(t,s;z)|\leq\exp(-\frac{\gamma}{2}|s-t|+|J|^{1-\tau/2}+2C_{0}|J|^{1-\tau})$
and
$|G^{\beta,\eta}_{J}(s,n^{\prime};z)|\leq\exp(-\frac{\gamma}{2}|s-n^{\prime}|+|J|^{1-\tau/2}+2C_{0}|J|^{1-\tau}).$
Therefore, we can get that
$|u(s)|$ $\leq 6\exp(-\frac{\gamma}{2}d+|J|^{1-\tau/2}+2C_{0}|J|^{1-\tau})$
$<6\exp(-\frac{\gamma}{2}d+C|J|^{1-\tau/2})<\exp(-\frac{\gamma}{12}d).$
For $s\in[n^{\prime\prime},n-1]$, the statement can be shown analogously. ∎
Lemma 3.28.
Assume $x_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $\tau,\nu$ be as in LDT and $l,n$ be integers such that $C(p,q,z_{0},\gamma)\leq l\leq n^{\nu/2}$. Assume that there exists an interval $I=[n^{\prime},n^{\prime\prime}]\subset[0,n-1]$ such that (3.27) holds and $|I|\geq l^{2/\nu}+\log n$. Then for any $(\omega,x)\in\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d}$ satisfying $|x-x_{0}|,|\omega-\omega_{0}|<\exp(-l)$ and any eigenvalue $|z_{j}^{[0,n-1]}(\omega,x)-z_{0}|<\exp(-l)/2$, we have
$|z_{j}^{[0,n-1]}(\omega,x)-z_{k}^{[0,n-1]}(\omega,x)|\geq\exp(-C|I|)$ (3.43)
for any $k\neq j$, where $C=C(\alpha_{i},z_{0})$.
Proof.
We argue by contradiction. Let $z_{1}=z_{j}^{[0,n-1]}(\omega,x)$, $z_{2}=z_{k}^{[0,n-1]}(\omega,x)$ and $u_{1}=u_{j}^{[0,n-1]}(\omega,x)$, $u_{2}=u_{k}^{[0,n-1]}(\omega,x)$ be the corresponding eigenfunctions and assume
$|z_{1}-z_{2}|<\exp(-C|I|).$
Since $u_{1},u_{2}$ are eigenfunctions corresponding to different eigenvalues, they are orthogonal and
$\|u_{1}-u_{2}\|^{2}=\|u_{1}\|^{2}+\|u_{2}\|^{2}.$
Let $\tilde{I}=\{s\in[0,n-1]:\mathrm{dist}(s,I)<l^{2/\nu}\}$. According to Lemma 3.27, one can obtain
$\sum_{s\notin\tilde{I}}|u_{i}(s)|^{2}\leq\|u_{i}\|^{2}\sum_{s\notin\tilde{I}}\exp(-\frac{\gamma}{12}\mathrm{dist}(s,I))\leq\|u_{i}\|^{2}\exp(-\frac{\gamma}{24}l^{2/\nu})$
(we used the fact that $l$ is taken to be large enough).
Therefore,
$\sum_{s\notin\tilde{I}}|u_{1}(s)-u_{2}(s)|^{2}\leq 2\exp(-\frac{\gamma}{24}l^{2/\nu})(\|u_{1}\|^{2}+\|u_{2}\|^{2}).$ (3.44)
For those $s\in\tilde{I}:=(I_{1},I_{2})$, applying Poisson formula, we have
$|u_{1}(s)-u_{2}(s)|$
$\leq 2\big{|}G_{[I_{1},I_{2}]}^{\beta,\eta}(I_{1},s;z_{1})-G_{[I_{1},I_{2}]}^{\beta,\eta}(I_{1},s;z_{2})\big{|}(|u_{1}(I_{1})|+|u_{1}(I_{1}+1)|+|u_{2}(I_{1})|+|u_{2}(I_{1}+1)|)$
$\quad+2\big{|}G_{[I_{1},I_{2}]}^{\beta,\eta}(s,I_{2};z_{1})-G_{[I_{1},I_{2}]}^{\beta,\eta}(s,I_{2};z_{2})\big{|}(|u_{1}(I_{2}-1)|+|u_{1}(I_{2})|+|u_{2}(I_{2})|+|u_{2}(I_{2}-1)|)$
$\leq 2(\|u_{1}\|+\|u_{2}\|)(\big{|}G_{[I_{1},I_{2}]}^{\beta,\eta}(I_{1},s;z_{1})-G_{[I_{1},I_{2}]}^{\beta,\eta}(I_{1},s;z_{2})\big{|}+\big{|}G_{[I_{1},I_{2}]}^{\beta,\eta}(s,I_{2};z_{1})-G_{[I_{1},I_{2}]}^{\beta,\eta}(s,I_{2};z_{2})\big{|}).$
Note that $|z_{j}^{[0,n-1]}(\omega,x)-z_{0}|<\exp(-l)/2$, then Lemma 3.25 can be applied. From (3.38), one can obtain that
$\big{|}G_{[I_{1},I_{2}]}^{\beta,\eta}(I_{1},s;z_{1})-G_{[I_{1},I_{2}]}^{\beta,\eta}(I_{1},s;z_{2})\big{|}+\big{|}G_{[I_{1},I_{2}]}^{\beta,\eta}(s,I_{2};z_{1})-G_{[I_{1},I_{2}]}^{\beta,\eta}(s,I_{2};z_{2})\big{|}<\frac{1}{4}.$
Thus, we have
$\sum_{s\in\tilde{I}}|u_{1}(s)-u_{2}(s)|^{2}<(\|u_{1}\|^{2}+\|u_{2}\|^{2}).$ (3.45)
Combining (3.44) and (3.45), we can obtain that
$\|u_{1}-u_{2}\|^{2}<\|u_{1}\|^{2}+\|u_{2}\|^{2},$
which contradicts the fact that $u_{1},u_{2}$ are orthogonal. ∎
4 Removing Double Resonances by Using Semialgebraic Sets
In this section, semialgebraic sets will be introduced by approximating the Verblunsky coefficients $\alpha_{i}$ with a polynomial $\tilde{\alpha}_{i}$. Then we obtain a result on elimination of double resonances using semialgebraic sets. More precisely, given $n\geq 1$, by truncating $\alpha_{i}$’s Fourier series and the Taylor series of the trigonometric functions, one can obtain a polynomial $\tilde{\alpha}_{i}$ of degree $\lesssim n^{4}$ such that
$\|\alpha_{i}-\tilde{\alpha}_{i}\|_{\infty}\lesssim\exp(-n^{2}).$ (4.1)
Let ~$\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}$ be the matrix with this truncated Verblunsky coefficient and the refined boundary condition $\beta,\eta\in\partial\mathbb{D}$. If we let $z_{j}^{[0,n-1]}(x,\omega)$ and $\tilde{z}_{j}^{[0,n-1]}(x,\omega)$ be the eigenvalues of $\mathcal{E}_{[0,n-1]}^{\beta,\eta}$ and ~$\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}$ respectively, then
$|z_{j}^{[0,n-1]}(x,\omega)-\tilde{z}_{j}^{[0,n-1]}(x,\omega)|\leq\|\mathcal{E}_{[0,n-1]}^{\beta,\eta}(x,\omega)-\tilde{\mathcal{E}}_{[0,n-1]}^{\beta,\eta}(x,\omega)\|\lesssim 8\|\alpha_{i}-\tilde{\alpha}_{i}\|_{\infty}.$ (4.2)
For the purpose of semialgebraic approximation, we take the following set into consideration
$\mathbb{T}_{n}^{d}(p,q)=\{\omega\in\mathbb{T}^{d}:\|k\cdot\omega\|\geq\frac{p}{|k|^{q}},\,\mathrm{for}\,\,\mathrm{all}\,\,\mathrm{nonzero}\,\,k\in\mathbb{Z}^{d},0<|k|\leq n\}.$
Definition 1.
[4, Definition 9.1] A set $\mathcal{S}\subset\mathbb{R}^{n}$ is called semialgebraic if it is a finite union of sets defined by a finite number of polynomial equalities and inequalities. More precisely, let $\mathcal{P}=\{P_{1},\ldots,P_{s}\}\subset R[X_{1},\ldots,X_{n}]$ be a family of real polynomials whose degrees are bounded by $d$. A (closed) semialgebraic set $\mathcal{S}$ is given by an expression
$\mathcal{S}=\bigcup\limits_{j}\bigcap\limits_{l\in\mathcal{L}_{j}}\{R^{n}|P_{l}s_{jl}0\},$ (4.3)
where $\mathcal{L}_{j}\subset\{1,\ldots,s\}$ and $s_{jl}\in\{\geq,\leq,=\}$ are arbitrary. We say that $\mathcal{S}$ has degree at most $sd$, and its degree is the infimum of $sd$ over all representations as in (4.3).
Lemma 4.1.
[4, Corollary 9.7] Let $\mathcal{S}\subset[0,1]^{n}$ be semialgebraic of degree $B$ and $\mathrm{mes}_{n}\mathcal{S}<\eta$. Let $\omega\in T^{n}$ satisfy a DC and N be a large integer,
$\log B\ll\log N<\log\frac{1}{\eta}.$
Then, for any $x_{0}\in T^{n}$
[…]
for some $\delta=\delta(\omega)$.
Lemma 4.2.
[3, Lemma 1.18] Let $\mathcal{A}\subset[0,1]^{s+r}$ be semialgebraic of degree $B$ and such that for each $t\in[0,1]^{r}$, $\mathrm{mes}_{s}(\mathcal{A}(t))<\eta$. Then
$\{(x_{1},\ldots,x_{2^{r}}):\mathcal{A}(x_{1})\cap\ldots\cap\mathcal{A}(x_{2^{r}})\neq\emptyset\}\subset[0,1]^{s^{2^{r}}}$
is semialgebraic of degree at most $B^{C}$ and measure at most
$\eta_{r}=B^{C}\eta^{s^{-r}2^{-r(r-1)/2}}\,\,with\,\,C=C(r).$
Note that in the above lemma $\mathcal{A}(t)=\{x\in[0,1]^{s}:(x,t)\in\mathcal{A}\}$ and $\mathcal{A}(x)=\{t\in[0,1]^{r}:(x,t)\in\mathcal{A}\}$.
Lemma 4.3.
[3, Lemma 1.20] Let $\mathcal{A}\subset[0,1]^{rs}$ be a semialgebraic set of degree $B$ and $\mathrm{mes}_{rs}(\mathcal{A})<\eta$. Let $\mathcal{N}_{1},\ldots,\mathcal{N}_{s-1}\subset\mathbb{Z}$ be finite sets with the property that
[…]
where $C=C(s,r)$. Assume also
$\max\limits_{n\in\mathcal{N}_{s-1}}|n|^{C}<\frac{1}{\eta}.$
Then
$\mathrm{mes}\{\omega\in[0,1]^{r}:(\omega,n_{1}\omega,\ldots,n_{s-1}\omega)\in\mathcal{A}\,\,for\,\,some\,\,n_{i}\in\mathcal{N}_{i}\}<B^{C}(\min\limits_{n\in\mathcal{N}_{1}}|n|)^{-1}.$
Lemma 4.4.
[4, Lemma 9.9] Let $\mathcal{S}\subset[0,1]^{d}$, $d=d_{1}+d_{2}$, be a semialgebraic set of degree $B$ and $\mathrm{mes}_{d}(\mathcal{S})<\vartheta$, $\log B\ll\log\frac{1}{\vartheta}$. We denote $(x,\omega)\in[0,1]^{d_{1}}\times[0,1]^{d_{2}}$ the product variable. Fix $\varepsilon>\vartheta^{\frac{1}{d}}$. Then there is a decomposition
$\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}$
$\mathcal{S}_{1}$ satisfying
$\mathrm{mes}_{d_{2}}(\mathrm{Proj}_{\omega}\mathcal{S}_{1})<B^{C}\varepsilon$
and $\mathcal{S}_{2}$ satisfying the transversality property
$\mathrm{mes}_{d_{1}}(\mathcal{S}_{2}\cap L)<B^{C}\varepsilon^{-1}\vartheta^{\frac{1}{d}}$
for any $d_{1}$-dimensional plane $L$ s.t. $\max\limits_{1\leq j\leq d_{2}}|\mathrm{Proj}_{L}(e_{j})|<\frac{1}{100}\varepsilon$ (we denote by $e_{1},\ldots,e_{d_{2}}$ the $\omega$-coordinate vectors).
From Corollary 3.12, we can get the following statements.
Corollary 4.5.
Assume $\omega\in\mathbb{T}^{d}(p,q)$, $z\in\partial\mathbb{D}$ and $L(\omega,z)>\gamma>0$. Then for any $x\in\mathbb{T}^{d}$ and $n\geq 1$ we have
$\|M_{n}(\omega,z;x)-\tilde{M}_{n}(\omega,z;x)\|\leq 8\|\alpha_{i}-\tilde{\alpha}_{i}\|_{\infty}\exp(nL(\omega,z)+Cn^{1-\tau}),$
where $C=C(p,q,z,\gamma)$. In particular,
$\big{|}\log\|M_{n}(\omega,z;x)\|-\log\|\tilde{M}_{n}(\omega,z;x)\|\big{|}\leq 8\|\alpha_{i}-\tilde{\alpha}_{i}\|_{\infty}\exp(nL(\omega,z)+Cn^{1-\tau}),$
$\big{|}\log|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|-\log|\tilde{\varphi}_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\big{|}\leq 8\|\alpha_{i}-\tilde{\alpha}_{i}\|_{\infty}\frac{\exp(nL(\omega,z)+Cn^{1-\tau})}{\max\{|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|,|\tilde{\varphi}_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\}},$
provided the right-hand sides are less than $\frac{1}{2}$.
Lemma 4.6.
Let $n\geq 1$, $\omega\in\mathbb{T}^{d}_{n}(p,q)$, $z\in\partial\mathbb{D}$, $\sigma,\tau$ be as in LDT, and
$\mathcal{B}_{n}:=\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x)\|-nL_{n}(\omega,z)|\geq 4n^{1-\tau}\}.$
Then there exists a semialgebraic set $\mathcal{S}_{n}$ such that $\mathcal{B}_{n}\subset\mathcal{S}_{n}$, $\mathrm{deg}(\mathcal{S}_{n})\leq n^{C}$, and $\mathrm{mes}(\mathcal{S}_{n})<\exp(-C_{0}n\sigma)$ provided $n\geq n(p,q,z)$.
Proof.
Take $\tilde{\alpha}_{n}$ as in (4.1) and let
$\mathcal{S}_{n}:=\{x\in\mathbb{T}^{d}:|\log\|\tilde{M}_{n}(\omega,z;x)\|_{\mathrm{HS}}-nL_{n}(\omega,z)|\geq 2n^{1-\tau}\},$
with $\|\cdot\|_{\mathrm{HS}}$ standing for the Hilbert-Schmidt norm. Obviously, $\mathrm{deg}(\mathcal{S}_{n})\leq n^{C}$. According to Corollary 4.5, $\mathcal{B}_{n}\subset\mathcal{S}_{n}$. In addition,
$\mathcal{S}_{n}\subset\{x\in\mathbb{T}^{d}:|\log\|M_{n}(\omega,z;x)\|-nL_{n}(\omega,z)|>n^{1-\tau}\}$
follows from Corollary 4.5.
Therefore, by LDT, the measure estimate of $\mathcal{S}_{n}$ follows. ∎
Lemma 4.7.
Let $\sigma,\tau$ be as in LDT. Then for any $n\geq n_{0}(p,q)$,
$C(p,q)\log n\leq\log J<n^{\sigma},$
$x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}_{J}(p,q)$, $z\in\partial\mathbb{D}$, we have
$\big{|}\frac{1}{J}\sum_{j=1}^{J}\log\|M_{n}(\omega,z;x+j\omega)\|-nL_{n}(\omega,z)\big{|}\leq 5n^{1-\tau}.$
Proof.
Let $\mathcal{B}_{n}$ and $\mathcal{S}_{n}$ be as in Lemma 4.7. For any $x\in\mathbb{T}^{d}\backslash\mathcal{S}_{n}$, we have that
$nL_{n}(\omega,z)-4n^{1-\tau}\leq\log\|M_{n}(\omega,z;x)\|\leq nL_{n}(\omega,z)+4n^{1-\tau}.$
While for $x\in\mathcal{S}_{n}$, one can obtain that
$0\leq\log\|M_{n}(\omega,z;x)\|\leq C(\alpha_{i},z)n.$
Select $J$ in accordance with the required assumptions. Combining Lemma 4.1, one can get
$\frac{1}{J}\sum_{j=1}^{J}\log\|M_{n}(\omega,z;x+j\omega)\|-nL_{n}(\omega,z)$
$\leq\frac{J-J^{1-\delta}}{J}(nL_{n}(\omega,z)+4n^{1-\tau})+\frac{J^{1-\delta}}{J}Cn-nL_{n}(\omega,z)$
$\leq\frac{J-J^{1-\delta}}{J}(4n^{1-\tau})+\frac{J^{1-\delta}}{J}Cn$
$\leq 5n^{1-\tau}.$
On the other hand,
$\frac{1}{J}\sum_{j=1}^{J}\log\|M_{n}(\omega,z;x+j\omega)\|-nL_{n}(\omega,z)$ $\geq\frac{J-J^{1-\delta}}{J}(nL_{n}(\omega,z)-4n^{1-\tau})-nL_{n}(\omega,z)$
$=\frac{J-J^{1-\delta}}{J}(-4n^{1-\tau})+\frac{J^{1-\delta}}{J}(-nL_{n}(\omega,z))$
$\geq-5n^{1-\tau}.$
Thus, we complete the proof. ∎
Lemma 4.8.
Let $n\geq 1$ and $\tau,\nu$ be as in LDT. Let $\mathcal{B}_{n}$ be the set of $(\omega,z,x)\in\mathbb{T}^{d}(p,q)\times\partial\mathbb{D}\times\mathbb{T}^{d}$ such that $L(\omega,z)>\gamma>0$ and
$\log|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\leq nL_{n}(\omega,z)-n^{1-\tau/2}.$
Then there exists a semialgebraic set $\mathcal{S}_{n}$ such that $\mathcal{B}_{n}\subset\mathcal{S}_{n}$, $\mathrm{deg}(\mathcal{S}_{n})\leq n^{C(p,q)}$, and $\mathrm{mes}(\mathcal{S}_{n})<\exp(-n^{\nu}/2)$, provided $n\geq N_{0}(\alpha_{i},p,q,\gamma)$. Furthermore,
$\mathrm{mes}(\mathcal{S}_{n}(\omega,z))<\exp(-n^{\nu}),\quad\mathcal{S}_{n}(\omega,z)=\{x:(\omega,z,x)\in\mathcal{S}_{n}\}.$
Proof.
Take $\tilde{\alpha}_{i}$ as in (4.1) and let $\mathcal{S}_{n}$ be the set of
$(\omega,z,x)\in\mathbb{T}_{J}^{d}(p,q)\times\partial\mathbb{D}\times\mathbb{T}^{d},\quad J=n^{C(a,b)}$
such that
$\frac{1}{nJ}\sum_{j=1}^{J}\log\|\tilde{M}_{n}(\omega,z;x+jw)\|_{HS}\geq\frac{\gamma}{2}$ (4.5)
and
$\log|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\leq\frac{1}{J}\sum_{j=1}^{J}\log\|\tilde{M}_{n}(\omega,z;x+jw)\|_{HS}-n^{1-\tau/2}/2.$ (4.6)
Obviously, $\mathrm{deg}(\mathcal{S}_{n})\leq n^{C}$. According to Corollary 4.5 and Lemma 4.6, $\mathcal{B}_{n}\subset\mathcal{S}_{n}$.
Combining Corollary 4.5, (4.5) and (4.6), we have $L_{n}(\omega,z)\geq\frac{\gamma}{4}$ and
$\log|\varphi_{[0,n-1]}^{\beta,\eta}(\omega,z;x)|\leq nL_{n}(\omega,z)-n^{1-\tau/2}/4.$
Therefore, $\mathrm{mes}(\mathcal{S}_{n}(\omega,z))<\exp(-n^{\nu})$ follows from LDT. In addition,
$\mathrm{mes}(\mathcal{S}_{n})<2\pi\exp(-n^{\nu})<\exp(-n^{\nu}/2).$
Lemma 4.9.
Let $l\geq 1$, $\gamma>0$, and $\tau,\nu$ be as in LDT. Let $\mathcal{N}_{1},\ldots,\mathcal{N}_{s-1}\subset\mathbb{Z}$, $s=2^{2d+1}$, be finite sets with the property that
[…]
and
$\max\limits_{n\in\mathcal{N}_{s-1}}|n|<\exp(cl^{\nu}),\,\,c=c(d).$
For any $l\geq l_{0}(\alpha_{i},p,q,\gamma)$ there exists a set $\Omega_{l}$, based on the selection of finite sets $\mathcal{N}_{i}$, such that
$\mathrm{mes}(\Omega_{l})\leq l^{C(p,q)}(\min\limits_{n\in\mathcal{N}_{1}}|n|)^{-1}$
and the subsequent statement is valid. For any $x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{l}$, $z\in\partial\mathbb{D}$, if $L(\omega,z)>\gamma$ and
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega,z;x)|\leq lL_{l}(\omega,z)-l^{1-\tau/2},$
then there exists $i\in\{1,\ldots,s-1\}$, depending on $\omega,z,x$, such that
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega,z;x+(n-1)\omega)|>lL_{l}(\omega,z)-l^{1-\tau/2},\,\,for\,\,all\,\,n\in\mathcal{N}_{i}.$
Proof.
Let $\mathcal{B}_{l}$, $\mathcal{S}_{l}$ be the sets from 4.8. Then $\mathcal{S}_{l}$ is semialgebraic, $\mathcal{B}_{l}\subset\mathcal{S}_{l}$, $\mathrm{deg}(\mathcal{S}_{l})\leq l^{C(p,q)}$, $\mathrm{mes}(\mathcal{S}_{l}(\omega,z))<\exp(-l^{\nu})$.
Let $\mathcal{T}_{l}$ be the set of $(\omega,z,x,y)\in\mathbb{T}^{d}_{l^{C}}(p,q)\times\partial\mathbb{D}\times\mathbb{T}^{d}\times\mathbb{T}^{d}$ such that $(\omega,z,x+y-\omega)\in\mathcal{S}_{l}$. Then $\mathcal{T}_{l}$ is a semialgebraic set and
$\mathrm{deg}(\mathcal{T}_{l})\leq l^{C},\,\,\mathrm{mes}(\mathcal{T}_{l}(\omega,z,x))<\exp(-l^{\nu}).$
According to Lemma 4.2,
$\mathcal{A}:=\{(y_{1},\ldots,y_{s}):\mathcal{T}_{l}(y_{1})\cap\cdots\cap\mathcal{T}_{l}(y_{s})\neq\emptyset\}$
is semialgebraic of degree at most $l^{C}$ and measure at most $\exp(-cl^{\nu})$, where $c=c(d)$.
Let $\Omega_{l}=\{\omega:(\omega,n_{1}\omega,\ldots,n_{s-1}\omega)\in\mathcal{A}\,\,\mathrm{for}\,\,\mathrm{some}\,\,n_{i}\in\mathcal{N}_{i}\}$. Then we can obtain the conclusion with the aid of Lemma 4.3. ∎
Corollary 4.10.
Let $l\geq 1$ and $\tau,\nu$ be as in LDT. Let $\mathcal{N}_{1},\ldots,\mathcal{N}_{(s-1)^{2}}\subset\mathbb{Z}$ be finite sets, with $s=2^{2d+1}$, possessing the property that
[…]
and
$\max\limits_{n\in\mathcal{N}_{(s-1)^{2}}}|n|<\exp(cl^{\nu}),\,\,c=c(d).$
For any $l\geq l_{0}(\alpha_{i},p,q,\gamma)$ there exists a set $\Omega_{l}$, based on the selection of finite sets $\mathcal{N}_{i}$, such that
$\mathrm{mes}(\Omega_{l})\leq l^{C(p,q)}(\min\limits_{n\in\mathcal{N}_{1}}|n|)^{-1}$
and the following statement holds. For any $x_{1},x_{2}\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{l}$, $z\in\partial\mathbb{D}$, if $L(\omega,z)>\gamma$ and
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega,z;x_{j})|\leq lL_{l}(\omega,z)-l^{1-\tau/2},\,\,j=1,2,$
then there exists $i\in\{1,\ldots,(s-1)^{2}\}$, depending on $\omega,z,x_{1},x_{2}$, such that
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega,z;x_{j}+(n-1)\omega)|>lL_{l}(\omega,z)-l^{1-\tau/2},\,\,for\,\,all\,\,n\in\mathcal{N}_{i}.$
Proof.
Let $\Omega_{l}$ be the union of the sets obtained from Lemma 4.9 with the following choices of finite sets:
$\mathcal{N}_{k(s-1)+1},\ldots,\mathcal{N}_{(k+1)(s-1)},\,\,k=0,\ldots,s-2,$ (4.7)
$\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{j},\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{(s-1)+j},\ldots,\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{(s-2)(s-1)+j}.$ (4.8)
It is evident that $\Omega_{l}$ complies with the prescribed measure bound. Let $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{l}$.
From Lemma 4.9 and (4.8), there exists $k\in\{0,\ldots,s-2\}$ such that
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega,z;x_{1}+(n-1)\omega)|>lL_{l}(\omega,z)-l^{1-\tau/2}$
for all $n\in\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{k(s-1)+j}$. Besides, from 4.9 and (4.7), there exists $j\in\{1,\ldots,s-1\}$ such that
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega,z;x_{2}+n\omega)|>lL_{l}(\omega,z)-l^{1-\tau/2}$
for all $n\in\mathop{\cup}\limits_{j=1}^{s-1}\mathcal{N}_{k(s-1)+j}$. Taking $i=k(s-1)+j$, the conclusion follows. ∎
For our purpose, we need the following Wegner’s estimate.
Lemma 4.11.
Let $\nu$ be as in LDT. Assume $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $l,n$ be integers such that $(2\log n)^{1/\nu}\leq Cl\leq n$. Then for any $n\geq N_{0}(p,q,\gamma)$ there exists a set $\mathcal{B}_{n,\omega_{0},z_{0}}$, $\mathrm{mes}(\mathcal{B}_{n,\omega_{0},z_{0}})<\exp(-Cl^{\nu}/2)$ such that for any $x\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega_{0},z_{0}}$ and any $(\omega,z)\in\mathbb{T}^{d}\times\partial\mathbb{D}$, $|x-x_{0}|$, $|\omega-\omega_{0}|$, $|z-z_{0}|<\exp(-l)$, we have
$\mathrm{dist}(z,\sigma(\mathcal{E}_{[0,n-1]}^{\beta,\eta}))\geq\exp(-l).$
Proof.
Take $\omega,z$ satisfying the assumptions. Let $\mathcal{B}_{n,\omega_{0},z_{0}}$ be the set of $x$ such that
$\log|\varphi^{\beta,\eta}_{[0,l-1]}(x+(m-1)\omega_{0},\omega_{0},z_{0})|<lL_{l}(\omega_{0},z_{0})-l^{1-\tau},\quad m\in[0,n-1].$
According to LDT, we can obtain
$\mathrm{mes}(\mathcal{B}_{n,\omega_{0},z_{0}})<\exp(-Cl^{v})<n\exp(-Cl^{v})<\exp(-\frac{Cl^{v}}{2}).$
By the covering form of LDT, for any $x\notin\mathcal{B}_{n,\omega_{0},z_{0}}$, we have
$\mathrm{dist}(z,\sigma(\mathcal{E}_{[0,n-1]}^{\beta,\eta}))\geq\exp(-l).$
Lemma 4.12.
Let $\varpi\in(0,1)$, $\bar{l}\geq 1$ and $\nu$ be as in LDT. Let $\Lambda_{0},\Lambda_{1}$ be intervals in $\mathbb{Z}$ such that $\bar{l}/10\leq|\Lambda_{0}|,|\Lambda_{1}|\leq 10\bar{l}$ and $0\in\Lambda_{0},\Lambda_{1}$. For any $x_{0}\in\mathbb{T}^{d}$, $\bar{l}\geq\bar{l_{0}}(\alpha_{i},p,q,\varpi,\gamma)$, $1\leq t\leq\exp(c_{0}\bar{l}^{\varpi\nu})$, there exists a set $\Omega_{\bar{l},t_{0},x_{0}}$, $\mathrm{mes}(\Omega_{\bar{l},t_{0},x_{0}})<\bar{l}^{C(p,q)}/t_{0}$ such that the following statement holds. For all $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{\bar{l},t_{0},x_{0}}$, $z\in\partial\mathbb{D}$, such that $L(\omega,z)>\gamma$, we have
$\min(\|(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0})-z)^{-1}\|,\|(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega,x_{0}+t\omega)-z)^{-1}\|)\leq\exp(\bar{l}^{\varpi})$
for all $t_{0}\leq|t|\leq\exp(c_{0}\bar{l}^{\varpi\nu})$.
Proof.
Fix $x_{0}\in\mathbb{T}^{d}$. Let $\mathcal{B}$ be the set of $(\omega,z,x)\in\mathbb{T}^{d}(p,q)\times\partial\mathbb{D}\times\mathbb{T}^{d}$, such that $L(\omega,z)>\gamma$,
$\|(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0})-z)^{-1}\|>\exp(\bar{l}^{\varpi}),\,\,\mathrm{and}\,\,\|(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega,x)-z)^{-1}\|>\exp(\bar{l}^{\varpi}).$
Let $\tilde{\alpha}_{i}$ be as in (4.1). Then $\mathcal{B}$ is contained in a semialgebraic set $\tilde{\mathcal{S}}$ of
$(\omega,z,x)\in\mathbb{T}_{J}^{d}(p,q)\times\partial\mathbb{D}\times\mathbb{T}^{d},\,\,J=\bar{l}^{C(p,q)}$
satisfying
$\frac{1}{|\Lambda_{1}|J}\sum_{j=1}^{J}\log\|\tilde{M}_{|\Lambda_{1}|}(\omega,z;x+j\omega)\|_{HS}\geq\frac{\gamma}{2},$
(~$\|(\tilde{\mathcal{E}}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0})-z)^{-1}\|>\frac{1}{2}\exp(\bar{l}^{\varpi}),\,\,\mathrm{and}\,\,\|(\tilde{\mathcal{E}}_{\Lambda_{1}}^{\beta,\eta}(\omega,x)-z)^{-1}\|>\frac{1}{2}\exp(\bar{l}^{\varpi}).$
It is clear that the degree of $\tilde{\mathcal{S}}$ is bounded above by $\bar{l}^{C}$. Furthermore, for $(\omega,z,x)\in\tilde{\mathcal{S}}$ one can obtain that
$L_{|\Lambda_{1}|}(\omega,z)\geq\frac{\gamma}{4},$
$\|(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0})-z)^{-1}\|>\frac{1}{4}\exp(\bar{l}^{\varpi}),\,\,\mathrm{and}\,\,\|(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega,x)-z)^{-1}\|>\frac{1}{4}\exp(\bar{l}^{\varpi}).$ (4.9)
For $(\omega,z,x)$ satisfying (4.9) and some $z_{0}\in\sigma(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega,x_{0}))$, we have $\mathrm{dist}(z_{0},\sigma(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega,x)))<8\exp(-\bar{l}^{\varpi})$. Taking a union over $z_{0}$ and according to Lemma 4.11, we have
$\mathrm{mes}(\mathrm{Proj}_{(\omega,x)}\tilde{\mathcal{S}})<\exp(-c\bar{l}^{\varpi\nu}).$
Set $\mathcal{S}:=\mathrm{Proj}_{(\omega,x)}\tilde{\mathcal{S}}.$ Let $\mathcal{S}=\mathcal{S}_{1}\cup\mathcal{S}_{2}$ be the decomposition of $\mathcal{S}$ corresponding to Lemma 4.4 with $\varepsilon=\frac{200}{t_{0}}$. In order to get the conclusion, we just need that $(\omega,\{x_{0}+t\omega\})\notin\mathcal{S}$.
Let $T$ be integers in the range $t_{0}\leq|t|\leq\exp(c_{0}\bar{l}^{\varpi\nu})$. Define
$\Omega_{\bar{l},t_{0},x_{0}}(\Lambda_{0},\Lambda_{1})=\mathrm{Proj}_{\omega}\mathcal{S}_{1}\cup(\mathop{\cup}\limits_{t\in T}\Omega_{\bar{l},t,x_{0}}(\Lambda_{0},\Lambda_{1})),$
$\Omega_{\bar{l},t,x_{0}}(\Lambda_{0},\Lambda_{1}):=\{\omega:(\omega,\{x_{0}+t\omega\})\in\mathcal{S}_{2}\},$
$\Omega_{\bar{l},t_{0},x_{0}}=\mathop{\cup}\limits_{\Lambda_{0},\Lambda_{1}}\Omega_{\bar{l},t_{0},x_{0}}(\Lambda_{0},\Lambda_{1}).$
Note that $\{\omega:(\omega,\{x_{0}+t\omega\})\in\mathcal{S}_{1}\}\subset\mathrm{Proj}_{\omega}\mathcal{S}_{1}$. From the assumptions, there are less than $C\bar{l}^{4}$ possible choices for $\Lambda_{0},\Lambda_{1}$. Then we only need to estimate the measure of $\Omega_{\bar{l},t_{0},x_{0}}$. Note that the set of $(\omega,\{x_{0}+t\omega\})$, $\omega\in[0,1]^{d}$, is contained in a union of planes $L_{i}$, $i\leq|t|^{d}$. The planes $L_{i}$ are parallel to $(\omega,t\omega\})$, $\omega\in\mathbb{R}^{d}$, then
$|\mathrm{Proj}_{L_{i}}e_{j}|\leq\frac{1}{|t|}\leq\frac{1}{t_{0}}<\frac{\varepsilon}{100}\,\,\mathrm{for}\,\,\mathrm{all}\,\,i,\,j,$
where $e_{j}$ are as in Lemma 4.4. By Lemma 4.4,
$\mathrm{mes}(\mathrm{Proj}_{\omega}\mathcal{S}_{1})<\bar{l}^{C}/t_{0},$
$\mathrm{mes}(\Omega_{\bar{l},t,x_{0}}(\Lambda_{0},\Lambda_{1}))=\sum_{i}\mathrm{mes}(\mathcal{S}_{2}\cap L_{i})\lesssim|t|^{d}\bar{l}^{C}t_{0}\exp(-c\bar{l}^{\varpi\nu})\leq\exp(-c^{\prime}\bar{l}^{\varpi\nu}),$
then we can obtain the conclusion. ∎
Corollary 4.13.
We use the notation and assumptions of Lemma 4.12. For any $\bar{l}\geq\bar{l}_{0}(\alpha_{i},p,q,\varpi,\gamma)$ and $1\leq t_{0}\leq\exp(c_{0}\bar{l}^{\varpi\nu})$, there exists a set $\Omega_{\bar{l},t_{0}}$, $\mathrm{mes}(\Omega_{\bar{l},t_{0}})<\bar{l}^{C(p,q)}t_{0}^{-1/2}$, such that for any $\omega\notin\Omega_{\bar{l},t_{0}}$ there exists a set $\mathcal{B}_{\bar{l},t_{0},\omega}$, $\mathrm{mes}(\mathcal{B}_{\bar{l},t_{0},\omega})<\bar{l}^{C(p,q)}t_{0}^{-1/2}$, and the following statement holds. For any $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{\bar{l},t_{0}}$, $x\in\mathbb{T}^{d}\backslash\mathcal{B}_{\bar{l},t_{0},\omega}$, $z\in\partial\mathbb{D}$,such that $L(\omega,z)>\gamma$, we have
$\min(\|(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega,x)-z)^{-1}\|,\|(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega,x+t\omega)-z)^{-1}\|)\leq\exp(\bar{l}^{\varpi})$
for all $t_{0}\leq|t|\leq\exp(c_{0}\bar{l}^{\varpi\nu})$.
Proof.
Let $\mathcal{B}_{\bar{l},t_{0}}$ be the set of $x\in\mathbb{T}^{d}$, $\omega\in\mathbb{T}^{d}(p,q)$ such that $\omega\in\Omega_{\bar{l},t_{0},x}$, with $\Omega_{\bar{l},t_{0},x}$ as in Lemma 4.12. With the aid of Chebyshev’s inequality, there exists a set $\Omega_{\bar{l},t_{0}}$, $\mathrm{mes}(\Omega_{\bar{l},t_{0}})\leq(\mathrm{mes}(\mathcal{B}_{\bar{l},t_{0}}))^{\frac{1}{2}}$ such that for $\omega\notin\Omega_{\bar{l},t_{0}}$ we have $\mathrm{mes}(\mathcal{B}_{\bar{l},t_{0},\omega})\leq(\mathrm{mes}(\mathcal{B}_{\bar{l},t_{0}}))^{\frac{1}{2}}$, where $\mathcal{B}_{\bar{l},t_{0},\omega}=\{x:(x,\omega)\in\mathcal{B}_{\bar{l},t_{0}}\}$. Then the conclusion follows by applying Lemma 4.12. ∎
5 Removing Double Resonances under NDR Condition
To begin with, we consider the local factorization for $\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)$ about the spectral variable, where $\Lambda$ is a finite interval in $\mathbb{Z}$. Inspired by [15, Section 4], we will introduce the theory about “no double resonances NDR” condition in this section.
Definition 2.
Let $\tau,\nu$ be as in LDT. An interval $\Lambda\subset\mathbb{Z}$ is said to be $(K,l,C)$-NDR with respect to $x_{0},\omega_{0},z_{0}$ if there exists a subset $\underline{\Lambda}\subset\Lambda$ with $|\underline{\Lambda}|\leq K$ such that
$\log|\varphi_{[0,l-1]}^{\beta,\eta}(\omega_{0},z_{0};x_{0}+(n-1)\omega_{0})|>lL_{l}(\omega_{0},z_{0})-Cl^{1-\tau/3}$
for all $n\in\Lambda\backslash\underline{\Lambda}$. Moreover, the connected components of $\Lambda\backslash\underline{\Lambda}$ must have length exceeding $l^{2/\nu}$. If $C=1$, $\Lambda$ is referred to as $(K,l)$-NDR.
For our purpose, we need the following Weierstrass’ Preparation Theorem and the perturbation theory of matrices.
Consider an analytic function $f(z,\omega_{1},\ldots,\omega_{d})$ defined in a polydisk
$\mathcal{P}_{*}=\mathcal{D}(z_{0},R_{0})\times\prod_{j=1}^{d}\mathcal{D}(\omega_{j,0},R_{0})\,\,z_{0},\,\omega_{j,0}\in\mathbb{C},\,\,R_{0}>0.$
Lemma 5.1.
[15, Lemma 2.28] Assume that $f(\cdot,\omega_{1},\ldots,\omega_{d})$ has no zeros on some circle
[…]
for any $\underline{\omega}=(\omega_{1},\ldots,\omega_{d})\in\mathcal{P}=\mathop{\prod}\limits_{j=1}^{d}\mathcal{D}(\omega_{j,0},r_{j,0})$ where $0<r_{j,0}<R_{0}$. Then there exist a polynomial $P(z,\underline{\omega})=z^{k}+a_{k-1}(\underline{\omega})z^{k-1}+\cdots+a_{0}(\underline{\omega})$ with $a_{j}(\underline{\omega})$ analytic in $\mathcal{P}$ and an analytic function $g(z,\underline{\omega})$, $(z,\underline{\omega})\in\mathcal{D}(z_{0},r)\times\mathcal{P}$ so that the following properties hold:
(a) $f(z,\underline{\omega})=P(z,\underline{\omega})g(z,\underline{\omega})$ for any $(z,\underline{\omega})\in\mathcal{D}(z_{0},r)\times\mathcal{P}$,
(b) $g(z,\underline{\omega})\neq 0$ for any $(z,\underline{\omega})\in\mathcal{D}(z_{0},r)\times\mathcal{P}$,
(c) for any $\underline{\omega}\in\mathcal{P}$, $P(\cdot,\underline{\omega})$ has no zeros in \$\mathbb{C}\backslash\mathcal{D}(z_{0},r)$.
Lemma 5.2.
[15, Lemma 2.37] Let $H,H_{0}$ be $n\times n$ matrices, $H_{0}$ is Hermitian, $E_{0}\in\mathbb{R}$, $r_{0}>0$. Assume the number of eigenvalues of $H_{0}$ in $(E_{0}-r_{0},E_{0}+r_{0})$ is at most $K$ and
$\|H-H_{0}\|\leq\frac{r_{0}}{32(K+1)^{2}}.$
Then there exists $r_{0}/2<r<r_{0}$, which depends only on $H_{0}$, such that $H$ and $H_{0}$ have the same number of eigenvalues in the disk $\mathcal{D}(E_{0},r)$. Moreover, neither $H$ nor $H_{0}$ have eigenvalues in the region
$r-\frac{r_{0}}{8(K+1)}\leq|\zeta-E_{0}|\leq r+\frac{r_{0}}{8(K+1)}.$
Remark 5.3.
In [15], during the proof of the above lemma, the authors used the basic property of Hermitian matrix $H_{0}$, that is, $\|(H_{0}-E)^{-1}\|=\frac{1}{\mathrm{dist}(E,\mathrm{spec}(H_{0}))}$ for $E\notin\mathrm{spec}(H_{0})$. Actually, this formula also holds true for unitary matrices. Therefore, we have the following corollary. Since the proof of the corollary below is similar to the above lemma, we will no longer prove it.
Corollary 5.4.
Let $\mathcal{E}_{0},\mathcal{E}_{1}$ be $n\times n$ matrices, $\mathcal{E}_{0}$ is unitary, $z_{0}\in\partial\mathbb{D}$, $r_{0}>0$. Assume the number of eigenvalues of $\mathcal{E}_{0}$ in $\mathcal{D}(z_{0},r_{0})\cap\partial\mathbb{D}$ is at most $K$ and
$\|\mathcal{E}_{1}-\mathcal{E}_{0}\|\leq\frac{r_{0}}{32(K+1)^{2}}.$
Then there exists $r_{0}/2<r<r_{0}$, which depends only on $\mathcal{E}_{0}$, such that $\mathcal{E}_{1}$ and $\mathcal{E}_{0}$ have the same number of eigenvalues in the disk $\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$. Moreover, neither $\mathcal{E}_{1}$ nor $\mathcal{E}_{0}$ have eigenvalues in the region
$\Big{\{}\zeta\in\partial\mathbb{D}:r-\frac{r_{0}}{8(K+1)}\leq|\zeta-z_{0}|\leq r+\frac{r_{0}}{8(K+1)}\Big{\}}.$
Lemma 5.5.
Suppose $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $K\geq 1$, $l\geq l_{0}(\alpha_{i},p,q,z_{0},\gamma)$, $r_{0}=\exp(-l)$. Assume $\Lambda$ is $(K,l)$-NDR with respect to $x_{0},\omega_{0},z_{0}$. There exists $\frac{r_{0}}{2}<r<r_{0}$ such that for any $(\omega,x)\in\mathbb{C}^{d}\times\mathbb{C}^{d}$ satisfying
$|x-x_{0}|<\frac{c(\alpha_{i})r_{0}}{(K+1)^{2}},\,\,|\omega-\omega_{0}|<\frac{c(\alpha_{i})r_{0}}{|\Lambda|(K+1)^{2}},$ (5.1)
we have
$\mathrm{dist}(\{z\in\partial\mathbb{D}:\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)=0\},\{z\in\partial\mathbb{D}:|z-z_{0}|=r\})\geq\frac{r_{0}}{8(K+1)}.$
Proof.
Due to (5.1),
$\|\mathcal{E}_{\Lambda}^{\beta,\eta}(\omega,x)-\mathcal{E}_{\Lambda}^{\beta,\eta}(\omega_{0},x_{0})\|\leq C(|x-x_{0}|+|\Lambda||\omega-\omega_{0}|)\leq\frac{r_{0}}{32(K+1)^{2}}.$
Then the conclusion follows from Corollary 5.4. ∎
Remark 5.6.
As we know, the restriction of Schrödinger operator to a finite interval is a Hermitian matrix which possesses many good properties. While the restriction of CMV matrix is a unitary matrix after modifying the boundary conditions whose properties are worse than those of a Hermitian matrix. For example, there is a famous Cauchy Interlacing Theorem for the eigenvalues of Hermitian matrices. But we can not expect such a result for unitary matrices. As a result, we can not estimate the number of the roots for $\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)=0$ in some disk. Consequently, when we study the factorization of $\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)$, we can not obtain the property of the factor of it as the property (d) in [15, Proposition 5.2]. However, without this property, we still can obtain the last statement of the following lemma.
Lemma 5.7.
Suppose $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma>0$. Let $\tau,\nu$ be as in LDT. Let $K\geq 1$, $l\geq l_{0}(\alpha_{i},p,q,z_{0},\gamma)$, $r_{0}=\exp(-l)$. Assume $\Lambda$ is $(K,l)$-NDR with respect to $x_{0},\omega_{0},z_{0}$. There exist
$P(\omega,x,z)=z^{k}+a_{k-1}(\omega,x)z^{k-1}+\cdots+a_{0}(\omega,x)$
with $a_{j}(\omega,x)$ analytic in the polydisk
$\mathcal{P}:=\{(\omega,x):|\omega-\omega_{0}|<cr_{0}|\Lambda|^{-1}(K+1)^{-2},|x-x_{0}|<cr_{0}(K+1)^{-2}\},$
and an analytic function $g(\omega,x,z)$ on $\mathcal{P}\times\mathcal{D}(z_{0},r)$, $r_{0}/2<r<r_{0}$ such that:
(a) $\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)=P(\omega,x,z)g(\omega,x,z)$ for any $(\omega,x,z)\in\mathcal{P}\times\mathcal{D}(z_{0},r)$;
(b) $g(\omega,x,z)\neq 0$ for any $(\omega,x,z)\in\mathcal{P}\times\mathcal{D}(z_{0},r)$;
(c) for any $(\omega,x)\in\mathcal{P}$, $P(\omega,x,\cdot)$ has no zeros in \$\mathbb{C}\backslash\mathcal{D}(z_{0},r)$;
(d) if $(\omega,x)\in\mathcal{P}\cap(\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d})$, $z\in\mathcal{D}(z_{0},r)$, and $\frac{r_{0}}{16(K+1)}\geq\exp(-|\Lambda|^{\nu/2})$,
then
$\log|g(\omega,x,z)|>|\Lambda|L_{|\Lambda|}(\omega,z)-|\Lambda|^{1-\tau/2}.$ (5.2)
Proof.
By Lemma 5.5, all the necessary conditions for Lemma 5.1 are satisfied, which leads to the conclusions (a)-(c). Then we need to verify (d).
Note that $P(\omega,x,z)$ is a product of factors $z-z_{j}^{\Lambda}(\omega,x)$ with $z_{j}^{\Lambda}(\omega,x)\in\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$, where $z_{j}^{\Lambda}(\omega,x)$ is the eigenvalue of $\mathcal{E}_{\Lambda}^{\beta,\eta}$. Thus, for $z\in\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$, $|P(\omega,x,z)|\leq(2r)^{k}<1$. Then we obtain
$\log|g(\omega,x,z)|>\log|\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)|$
for any $z\in\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$.
Let $r-\frac{r_{0}}{16(K+1)}<r^{\prime}<r$. From Lemma 5.5, for any $z\in\{z\in\partial\mathbb{D}:|z-z_{0}|=r^{\prime}\}$ we have
$\mathrm{dist}(z,\sigma(\mathcal{E}_{\Lambda}^{\beta,\eta}(\omega,x)))\geq\frac{r_{0}}{16(K+1)}\geq\exp(-|\Lambda|^{\nu/2}).$
According to the spectral form of LDT, we have
$\log|\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)|>|\Lambda|L_{|\Lambda|}(\omega,z)-|\Lambda|^{1-\tau/2}.$
Therefore, (5.2) holds. Since the left-hand side of (5.2) is harmonic (in $z$) and the right-hand side is subharmonic, it follows that the estimate holds for all $z\in\mathcal{D}(z_{0},r^{\prime})\cap\partial\mathbb{D}$. Since this is true for $r^{\prime}$ arbitrarily close to $r$, the conclusion follows. ∎
Let us recall some basic theory about the resultant of two polynomials that we will use later.
Definition 3.
[15, Section 2.9] Let $f(z)=a_{k}z^{k}+a_{k-1}z^{k-1}+\cdots+a_{0}$, $g(z)=b_{m}z^{m}+b_{m-1}z^{m-1}+\cdots+b_{0}$ be polynomials with complex coefficients. Let $\zeta_{i}$, $1\leq i\leq k$ and $\eta_{j}$, $1\leq j\leq m$ be the zeros of $f(z)$ and $g(z)$ respectively. The resultant of $f$ and $g$ is defined as follows:
$\mathrm{Res}(f,g)=a_{k}^{m}b_{m}^{k}\prod_{i,j}(\zeta_{i}-\eta_{j}).$ (5.3)
Lemma 5.8.
[15, Lemma 2.29] Let $f$, $g$ be polynomials as above and set
$s=\max(k,m),\,\,r=\max(\max|\zeta_{i}|,\max|\eta_{j}|).$
Let $\delta\in(0,1)$. If $|\mathrm{Res}(f,g)|>\delta$ and $r\leq\frac{1}{2}$, then
$\max(|f(z)|,|g(z)|)>(\delta/2)^{s}\,\,for\,\,all\,\,z.$
Lemma 5.9.
Suppose $x_{0}\in\mathbb{T}^{d}$, $\omega_{0}\in\mathbb{T}^{d}(p,q)$, $z_{0}\in\partial\mathbb{D}$ and $L(\omega_{0},z_{0})>\gamma$. Let $\tau,\nu$ be as in LDT and $1\leq l\leq\bar{l}\leq\exp(l)$, $1\leq K\leq\bar{l}^{\frac{1-\tau}{10}}$, $\exp(l^{2})\leq|t|\leq\exp(\bar{l}^{\frac{\nu(1-\tau)}{4}})$. Let $\Lambda_{j}$, $j=0,1$, be intervals in $\mathbb{Z}$ such that $\bar{l}/10\leq|\Lambda_{j}|\leq 10\bar{l}$, $0\in\Lambda_{j}$, and $\Lambda_{j}$ is $(K,l)$-NDR with respect to $x_{j}$, $\omega_{0}$, $z_{0}$, $x_{j}:=x_{0}+jt\omega_{0}$, $j=0,1$. For any $l\geq l_{0}(\alpha_{i},p,q,z_{0},\gamma)$ there exists a set $\Omega_{\bar{l}}$, $\mathrm{mes}(\Omega_{\bar{l}})<\exp(-l^{2}/4)$, such that if $\omega_{0}\notin\Omega_{\bar{l}}$ the following statement holds. There exists a set $\mathcal{B}_{\bar{l},t}(x_{0},\omega_{0},z_{0})$, such that
$\mathrm{mes}(\mathcal{B}_{\bar{l},t})<\exp(-\bar{l}^{\frac{1-\tau}{6d}})/|t|^{d},$
and for any $(\omega,x)\in(\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d})\backslash\mathcal{B}_{\bar{l},t}$, $z\in\partial\mathbb{D}$,
$|x-x_{0}|<\exp(-4l),\,\,|\omega-\omega_{0}|<\exp(-4l)/|t|,\,\,|z-z_{0}|<\exp(-l)/2,$ (5.4)
we have
$\max_{j=0,1}(\log|\varphi_{\Lambda_{j}}^{\beta,\eta}(\omega,z;x+jt\omega)|-|\Lambda_{j}|L_{|\Lambda_{j}|}(\omega,z)+2|\Lambda_{j}|^{1-\tau/2})>0.$
Proof.
According to Lemma 3.7, $L(\omega_{0},z)>\frac{\gamma}{2}$ for any $|z-z_{0}|<\exp(-l)$ (provided $l$ is large enough). Let $\Omega_{\bar{l},t_{0}}$ $\mathcal{B}_{\bar{l},t_{0},\omega_{0}}$ be the sets from Corollary 4.13, then
$\mathrm{mes}(\Omega_{\bar{l},t_{0}}),\mathrm{mes}(\mathcal{B}_{\bar{l},t_{0},\omega_{0}})<\exp(-l^{2}/4).$
Set $\Omega_{\bar{l}}:=\Omega_{\bar{l},t_{0}}$. Suppose $\omega_{0}\notin\Omega_{\bar{l}}$. Let $x_{0}^{\prime}\in\mathbb{T}^{d}$, $|x_{0}^{\prime}-x_{0}|<\exp(-l^{2}/(5d))$ be such that $x_{0}^{\prime}\notin\mathcal{B}_{\bar{l},t_{0},\omega_{0}}$. We can obtain
$\min(\|(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega_{0},x_{0}^{\prime})-z)^{-1}\|,\|(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega_{0},x_{0}^{\prime}+t\omega_{0})-z)^{-1}\|)\leq\exp(\bar{l}^{\frac{1-\tau}{3}})$ (5.5)
for all $|z-z_{0}|<\exp(-l)$ and $\exp(l^{2})\leq|t|\leq\exp(c\bar{l}^{\nu(1-\tau)/3})$.
Let $\varphi_{\Lambda_{j}}^{\beta,\eta}(\omega,z;x)=P_{j}(\omega,x,z)g_{j}(\omega,x,z)$, $j=0,1$, be the factorizations from Lemma 5.7. We know that
$P_{j}(\omega,x,z)=z^{k_{j}}+a_{j,k_{j}-1}(\omega,x)z^{k_{j}-1}+\cdots+a_{j,0}(\omega,x)$
with $a_{j,i}(\omega,x)$ analytic in a polydisk containing
$\mathcal{P}_{j}:=\{(\omega,x)\in\mathbb{C}^{2d}:|\omega-\omega_{0}|<\exp(-3l),|x-x_{j}|<\exp(-2l)\},$
and all the zeros of $P_{j}(\omega,x,\cdot)$, $(\omega,x)\in\mathcal{P}_{j}$, are contained in $\mathcal{D}(z_{0},r)\cap\partial\mathbb{D}$.
Due to $\exp(l^{2})\leq|t|\leq\exp(\bar{l}^{\frac{\nu(1-\tau)}{4}})$, then $\bar{l}\geq l^{\frac{8}{\nu(1-\tau)}}$. In addition,
$\frac{r_{0}}{16(K+1)}=\frac{\exp(-l)}{16(K+1)}\geq\frac{\exp(-|\Lambda_{j}|^{\frac{\nu(1-\tau)}{8}})}{16(K+1)}\geq\exp(-|\Lambda_{j}|^{\nu/2}).$
Thus, the condition needed for (d) of Lemma 5.7 is satisfied. Then we have
$\log|g_{j}(\omega,x,z)|>|\Lambda_{j}|L_{|\Lambda_{j}|}(\omega,z)-|\Lambda_{j}|^{1-\tau/2}$ (5.6)
for any $(\omega,x)\in\mathcal{P}_{j}\cap(\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d})$ and $z\in\mathcal{D}(z_{0},\exp(-l)/2)\cap\partial\mathbb{D}$.
Let $R(\omega,x)=\mathrm{Res}(P_{0}(\omega,x,\cdot),P_{1}(\omega,x+t\omega,\cdot))$. Then $R$ is defined on the polydisk
$\mathcal{P}:=\{(\omega,x)\in\mathbb{C}^{2d}:|\omega-\omega_{0}|<\exp(-3l)/|t|,|x-x_{j}|<\exp(-3l)\}.$
According to the definition of the resultant and the properties of $P_{j}$, we get
$R(\omega,x)=\prod(z_{i}^{\Lambda_{0}}(\omega,x)-z_{j}^{\Lambda_{1}}(\omega,x+t\omega)),$
where the product is only in terms of eigenvalues contained in $\mathcal{D}(z_{0},\exp(-l))\cap\partial\mathbb{D}$. Then we have that $\sup|R(\omega,x)|\leq(2\exp(-l))^{k_{0}k_{1}}<1$.
From (5.5), for $z\in(\sigma(\mathcal{E}_{\Lambda_{0}}^{\beta,\eta}(\omega_{0},x_{0}^{\prime}))\cup\sigma(\mathcal{E}_{\Lambda_{1}}^{\beta,\eta}(\omega_{0},x_{0}^{\prime}+t\omega_{0})))\cap\mathcal{D}(z_{0},\exp(-l))\cap\partial\mathbb{D}$ we have that
$|z_{i}^{\Lambda_{0}}(\omega_{0},x_{0})-z_{j}^{\Lambda_{1}}(\omega_{0},x_{0}^{\prime}+t\omega_{0})|\geq\exp(-\bar{l}^{\frac{1-\tau}{3}})$
for the pairs of eigenvalues contained in $\mathcal{D}(z_{0},\exp(-l))\cap\partial\mathbb{D}$. Therefore,
$|R(\omega_{0},x_{0}^{\prime})|\geq(\exp(-\bar{l}^{\frac{1-\tau}{3}}))^{k_{0}k_{1}}\geq\exp(-|\Lambda_{0}||\Lambda_{0}|\bar{l}^{\frac{1-\tau}{3}})\geq\exp(-100\bar{l}^{2}\bar{l}^{\frac{1-\tau}{3}}).$
Applying Cartan’s estimate to $R(\omega,x)$ with $M=0$, $m=-100\bar{l}^{2}\bar{l}^{\frac{1-\tau}{3}}$, $H=\bar{l}^{\frac{1-\tau}{3}}$, one can obtain that
$\log|R(\omega,x)|>-100C\bar{l}^{2}\bar{l}^{\frac{1-\tau}{3}},$
which implies that $|R(\omega,x)|>\exp(-100C\bar{l}^{2}\bar{l}^{\frac{1-\tau}{3}})$ for all $(\omega,x)\in(\frac{1}{6}\mathcal{P}\cap\mathbb{R}^{2d})\backslash\mathcal{B}$, where $\mathcal{B}=\mathcal{B}(\Lambda_{0},\Lambda_{1})$ and
$\mathrm{mes}(\mathcal{B})\leq\frac{C\exp(-6dl)}{|t|^{d}}\exp(-H^{\frac{1}{2d}}).$
Note that $\frac{1}{6}\mathcal{P}\cap\mathbb{R}^{2d}$ contains the points $(\omega,x)$ satisfying (5.4). Let $\mathcal{B}_{\bar{l},t}=\cup\mathcal{B}(\Lambda_{0},\Lambda_{1})$. According to Lemma 5.3, one can obtain
$\max(|P_{0}(\omega,x,z)|,|P_{1}(\omega,x+t\omega,z)|)\geq(\frac{1}{2}\exp(-100C\bar{l}^{2}\bar{l}^{\frac{1-\tau}{3}}))^{10\bar{l}}>\exp(-\bar{l}^{5})$
for all $z$ and $(\omega,x)\notin\mathcal{B}_{\bar{l},t}$. Therefore, the conclusion follows by applying the factorizations of the determinants and (5.6). ∎
Lemma 5.10.
Let $\tau,\nu$ be as in LDT, $1\leq l\leq\bar{l}\leq\exp(l)$, $1\leq K\leq\bar{l}^{\frac{1-\tau}{10}}$, and $\exp(l^{2})\leq n$. Let $\Lambda_{j}$, $j=0,1$, be intervals in $\mathbb{Z}$ such that $\bar{l}/10\leq|\Lambda_{j}|\leq 10\bar{l}$, $0\in\Lambda_{j}$, $j=0,1$. For $l\geq l_{0}(\alpha_{i},p,q,\gamma)$ there exists $\varrho=\varrho(p,q)$ such that if $n\leq\exp(\bar{l}^{\varrho})$, there exists a set $\Omega_{n,\bar{l}}$, $\mathrm{mes}(\Omega_{n,\bar{l}})<2\exp(-l^{2}/4)$, and for every $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n,\bar{l}}$ there exists a set $\mathcal{B}_{n,\bar{l},\omega}$, $\mathrm{mes}(\mathcal{B}_{n,\bar{l},\omega})<\exp(-\bar{l}^{\varrho})$ such that the following statement holds. For any $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n,\bar{l}}$, $x\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\bar{l},\omega}$, $z\in\partial\mathbb{D}$, and $m_{0},m_{1}\in[0,n-1]$ such that $|m_{0}-m_{1}|\geq\exp(l^{2})$, we have that if $L(\omega,z)>\gamma$ and $\Lambda_{j}$, $j=0,1$, are $(K,l,\frac{1}{2})$-NDR intervals with respect to $x+m_{j}\omega$, $\omega$, $z$, then
$\max_{j=0,1}(\log|\varphi_{\Lambda_{j}}^{\beta,\eta}(\omega,z;x+m_{j}\omega)|-|\Lambda_{j}|L_{|\Lambda_{j}|}(\omega,z)+2|\Lambda_{j}|^{1-\tau/2})>0.$
Proof.
Let $\Omega_{\bar{l}}$ be the set from Lemma 5.9 (with $\frac{\gamma}{2}$ instead of $\gamma$). Let
$\{x:|x-x_{k}|<\exp(-l^{3/2})\},\,\,k\lesssim\exp(dl^{3/2})$
$\{\omega:|\omega-\omega_{k^{\prime}}|<\exp(-l^{3/2})/n\},\,\,k\lesssim n^{d}\exp(dl^{3/2})$
$\{z:|z-z_{k^{\prime\prime}}|<\exp(-l^{3/2})\},\,\,k^{\prime\prime}\lesssim\exp(l^{3/2})$
be covers of $\mathbb{T}^{d}$, $\mathbb{T}^{d}\backslash\Omega_{\bar{l}}$ and $\partial\mathbb{D}$, respectively. Note that if $|\omega-\omega_{k^{\prime}}|<\exp(-l^{3/2})/n$, $|z-z_{k^{\prime\prime}}|<\exp(-l^{3/2})$, and $L(\omega,z)>\gamma$, then $L(\omega_{k^{\prime}},z_{k^{\prime\prime}})>\frac{\gamma}{2}$ (provided $l$ is large enough) by using Lemma 3.7. If $L(\omega_{k^{\prime}},z_{k^{\prime\prime}})>\frac{\gamma}{2}$, we let $\mathcal{B}_{\bar{l},t}(x_{k},\omega_{k^{\prime}},z_{k^{\prime\prime}})$ be the set from Lemma 5.9 (with $\frac{\gamma}{2}$ instead of $\gamma$). Otherwise we let $\mathcal{B}_{\bar{l},t}(x_{k},\omega_{k^{\prime}},z_{k^{\prime\prime}})$ be the empty set.
Let $\mathcal{B}_{n,\bar{l}}=\mathop{\cup}\limits_{k,k^{\prime},k^{\prime\prime},t,m}S_{m}(\mathcal{B}_{\bar{l},t}(x_{k},\omega_{k^{\prime}},z_{k^{\prime\prime}}))$, where $S_{m}(x,\omega)=(x-m\omega,\omega)$, $m\in[0,n-1]$, $|t|\in[\exp(l^{2}),n]$. Then
$\mathrm{mes}(\mathcal{B}_{n,\bar{l}})\leq Cc^{d+2}\exp((2d+1)l^{3/2})\exp(-\bar{l}^{\frac{\nu(1-\tau)}{6d}}).$
From the restriction of $|t|$ in Lemma 5.9, we require that $n\leq\exp(\bar{l}^{\frac{\nu(1-\tau)}{4}})$. Take $\varrho=\varrho(\tau,\nu)=\varrho(p,q)$ such that if $n\leq\exp(\bar{l}^{\varrho})$, the above restriction is satisfied and
$\mathrm{mes}(\mathcal{B}_{n,\bar{l}})<\exp(-2\bar{l}^{\varrho}).$
According to Chebyshev’s inequality, there exists a set $\tilde{\Omega}_{n,\bar{l}}$, $\mathrm{mes}(\tilde{\Omega}_{n,\bar{l}})<\exp(-\bar{l}^{\varrho})$ such that for $\omega\notin\tilde{\Omega}_{n,\bar{l}}$ we have $\mathrm{mes}(\mathcal{B}_{n,\bar{l},\omega})<\exp(-\bar{l}^{\varrho})$, where $\mathcal{B}_{n,\bar{l},\omega}=\{x:(x,\omega)\in\mathcal{B}_{n,\bar{l}}\}$.
Set $\Omega_{n,\bar{l}}=\Omega_{\bar{l}}\cup\tilde{\Omega}_{n,\bar{l}}$. Let $x,\omega,z,m_{j},\Lambda_{j}$ be as in the assumptions. Then there exist $x_{k},\omega_{k^{\prime}},z_{k^{\prime\prime}}$ such that
$|x+m_{0}\omega-x_{k}|<\exp(-l^{3/2}),\,\,|\omega-\omega_{k^{\prime}}|<\exp(-l^{3/2})/n,\,\,|z-z_{k^{\prime\prime}}|<\exp(-l^{3/2}).$
Due to $L(\omega,z)>\gamma$, then we have $L(\omega_{k^{\prime}},z_{k^{\prime\prime}})>\frac{\gamma}{2}$. Since $\Lambda_{j}$, $j=0,1$, are $(K,l,\frac{1}{2})$-NDR intervals with respect to $x+m_{j}\omega,\omega,z$, then $\Lambda_{j}$, $j=0,1$, are $(K,l)$-NDR intervals with respect to $x_{k}+jt\omega_{k^{\prime}},\omega_{k^{\prime}},z_{k^{\prime\prime}},t=m_{1}-m_{0}$.
The choice of above exceptional sets implies that $(x+,m_{0}\omega,\omega)\notin\mathcal{B}_{\bar{l},t}(x_{k},\omega_{k^{\prime}},z_{k^{\prime\prime}})$. Then according to Lemma 5.9, we have
$\max_{j=0,1}(\log|\varphi_{\Lambda_{j}}^{\beta,\eta}(\omega,z;x+m_{0}\omega+jt\omega)|-|\Lambda_{j}|L_{|\Lambda_{j}|}(\omega,z)+2|\Lambda_{j}|^{1-\tau/2})>0.$
Due to $t=m_{1}-m_{0}$, then $x+m_{0}\omega+jt\omega=x+m_{0}\omega+j(m_{1}-m_{0})\omega=x+m_{j}\omega$, $j=0,1$.
Thus, the conclusion follows. ∎
6 Full Scale Localization
In this section, we formulate the full scale localization. To do that, we require the following lemmas.
Lemma 6.1.
[15, Section 2.40] Let $A$ be an $N\times N$ Hermitian matrix. Let $E,\varepsilon\in\mathbb{R}$, $\varepsilon>0$ and suppose there exists $\phi\in\mathbb{R}^{N}$, $\|\phi\|=1$, such that
$\|(A-E)\phi\|<\varepsilon.$
Then the following statements hold:
(a) There exists a normalized eigenvector $\psi$ of $A$ with an eigenvalue $E_{0}$ such that
$E_{0}\in(E-\varepsilon\sqrt{2},E+\varepsilon\sqrt{2}),$
$|\langle\phi,\psi\rangle|\geq(2N)^{-1/2}.$
(b) If in addition there exists $\hat{\varepsilon}>\varepsilon$ such that the subspace of the eigenvectors of $A$ with eigenvalues falling into the interval $(E-\hat{\varepsilon},E+\hat{\varepsilon})$ is at most of dimension one, then there exists a normalized eigenvector $\psi$ of $A$ with an eigenvalue $E_{0}\in(E-\varepsilon,E+\varepsilon)$, such that
$\|\phi-\psi\|<\sqrt{2}\hat{\varepsilon}^{-1}\varepsilon.$
From the above lemma and its proof, it is obvious that this Lemma can also be applicable to a unitary matrix. Then we have the following corollary. Since the proof is similar to the above lemma, we no longer prove it.
Corollary 6.2.
Let $A$ be an $N\times N$ unitary matrix. Let $z\in\partial\mathbb{D}$, $\tilde{\varepsilon}>0\in\mathbb{R}$, and suppose there exists $\phi\in\mathbb{C}^{N}$, $\|\phi\|=1$, such that
$\|(A-z)\phi\|<\tilde{\varepsilon}.$
Then the following statements hold:
(a) There exists a normalized eigenvector $\psi$ of $A$ with an eigenvalue $z_{0}$ such that
$z_{0}\in\mathcal{D}(z,\tilde{\varepsilon}\sqrt{2})\cap\partial\mathbb{D},$
$|\langle\phi,\psi\rangle|\geq(2N)^{-1/2};$
(b) If in addition there exists $\hat{\varepsilon}>\tilde{\varepsilon}$ such that the subspace of the eigenvectors of $A$ with eigenvalues falling into the interval $\mathcal{D}(z,\hat{\varepsilon})\cap\partial\mathbb{D}$ is at most of dimension one, then there exists a normalized eigenvector $\psi$ of $A$ with an eigenvalue $z_{0}\in\mathcal{D}(z,\tilde{\varepsilon})\cap\partial\mathbb{D}$, such that
$\|\phi-\psi\|<\sqrt{2}\hat{\varepsilon}^{-1}\tilde{\varepsilon}.$
Lemma 6.3.
Let $\varepsilon\in(0,\frac{1}{5})$ and $\tau,\nu$ be as in LDT. For $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, there exists a set $\Omega_{n}$,
$\mathrm{mes}(\Omega_{n})<\exp(-(\log n)^{\varepsilon\nu}),$
such that for $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$ there exists a set $\mathcal{B}_{n,\omega}$,
$\mathrm{mes}(\mathcal{B}_{n,\omega})<\exp(-(\exp(\log n)^{\varepsilon\nu})^{\varrho})$
and the following holds for any $\omega\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$, $x\in\mathbb{T}^{d}(p,q)\backslash\mathcal{B}_{n,\omega}$ and $z\in\partial\mathbb{D}$ such that $L(\omega,z)>\gamma$, where $\varrho$ is as in Lemma 5.10. Let $1\ll\underline{C}_{1}\ll\overline{C}_{1}\ll\cdots\ll\underline{C}_{(s-1)^{2}}\ll\overline{C}_{(s-1)^{2}}$, $s=2^{2d+1}$, be constants such that the sets
$\mathcal{N}_{k}=\{N\in\mathbb{Z}:\exp(\underline{C}_{k}l^{\nu/2})\leq|N|\leq\exp(\overline{C}_{k}l^{\nu/2})\}$
satisfy the assumptions of Corollary 4.10, where $l=\lceil(\log n)^{2\varepsilon}\rceil$.
If $m_{0}\in[0,n-1]$ is such that
$\log|\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x)|\leq|\Lambda|L_{|\Lambda|}(\omega,z)-|\Lambda|^{1-\tau/4}$
for all intervals $\Lambda\subset[0,n-1]$ satisfying $\mathrm{dist}(m_{0},[0,n-1]\backslash\Lambda)>|\Lambda|/100$, $|\Lambda|\geq(\log n)^{\varepsilon}$, then for any $\bar{l}\geq(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}$ we have
$\log|\varphi_{\Lambda}^{\beta,\eta}(\omega,z;x+(m-1)\omega)|>\bar{l}L_{\bar{l}}(\omega,z)-|\Lambda|^{1-\tau/2}$ (6.1)
for any $m\in[0,n-\bar{l}]$ such that $|m-m_{0}|\geq\bar{l}+2\exp(l^{2})$.
Proof.
Let $\bar{l}_{k}=\lfloor\exp(\overline{C}_{k}l^{\nu/2})\rfloor$, $K_{k}=\lceil\exp(\underline{C}_{k}l^{\nu/2})\rceil$. Taking the constants $\overline{C}_{k},\underline{C}_{k}$ such that $K_{k}\leq\bar{l}_{k}^{\frac{1-\tau}{10}}$, then the condition needed for Lemma 5.10 is satisfied. We choose
$\Omega_{n}:=\Omega_{l}\cup(\mathop{\cup}\limits_{k}\Omega_{n,\bar{l}_{k}}),\,\,\mathcal{B}_{n,\omega}:=\mathop{\cup}\limits_{k}\mathcal{B}_{n,\bar{l}_{k},\omega},$
where $\Omega_{l}$ is the set from Corollary 4.10 and $\Omega_{n,\bar{l}_{k}}$, $\mathcal{B}_{n,\bar{l}_{k},\omega}$ are the sets from Lemma 5.10. Then one can obtain that
$\mathrm{mes}(\Omega_{n})<l^{C(p,q)}\exp(-\underline{C}_{1}l^{\nu/2})+2(s-1)^{2}\exp(-l^{2}/4)<\exp(-l^{\nu/2})<\exp(-(\log n)^{\varepsilon\nu}),$
$\mathrm{mes}(\mathcal{B}_{n,\omega})<(s-1)^{2}\exp(-\bar{l}_{k}^{\varrho})<(s-1)^{2}\exp(-\bar{l}_{1}^{\varrho})<\exp(-(\exp(\log n)^{\varepsilon\nu})\varrho).$
Let $\Lambda_{0}=[m_{0}^{\prime},m_{0}^{\prime\prime}]\subset[0,n-1]$, $|\Lambda_{0}|=l$, be an interval such that $\mathrm{dist}(m_{0},[0,n-1]\backslash\Lambda_{0})>|\Lambda_{0}|/100$. Then this implies that
$\log|\varphi_{\Lambda_{0}}^{\beta,\eta}(\omega,z;x)|=\log|\varphi_{l}^{\beta,\eta}(\omega,z;x+(m_{0}^{\prime}-1)\omega)|>lL_{l}(\omega,z)-l^{1-\tau/2}.$
According to Corollary 4.10, for each $m^{\prime}\in[0,n-1]$ we either have
$\log|\varphi_{l}^{\beta,\eta}(\omega,z;x+(m^{\prime}-1)\omega)|>lL_{l}(\omega,z)-l^{1-\tau/2}$
(there exists $k\in\{1,\ldots,(s-1)^{2}\}$ such that the above inequality holds for all $m^{\prime}\in\mathcal{N}_{k}$) or there exists $k=k(m_{0}^{\prime},m^{\prime})$ such that $[-\bar{l}_{k},\bar{l}_{k}]$ is $(K_{k},l,\frac{1}{2})$-NDR with respect to both $x+(m_{0}^{\prime}-1)\omega,\omega,z$ and $x+(m^{\prime}-1)\omega,\omega,z$.
From Lemma 5.10, for any $m^{\prime}\in[0,n-1]$ such that $|m^{\prime}-m_{0}^{\prime}|\geq\exp(l^{2})$, we have
$\log|\varphi_{\Lambda(m^{\prime})}^{\beta,\eta}(\omega,z;x+(m^{\prime}-1)\omega)|>|\Lambda(m^{\prime})|L_{|\Lambda(m^{\prime})|}(\omega,z)-2|\Lambda(m^{\prime})|^{1-\tau/2},$ (6.2)
where $\Lambda(m^{\prime})$ is either $[0,l-1]$ or $[-\bar{l}_{k},\bar{l}_{k}]$ with $k=k(m_{0}^{\prime},m^{\prime})$ as above.
According to the definition of NDR intervals, if $[-\bar{l}_{k},\bar{l}_{k}]$ is NDR with respect to $x+(m_{0}^{\prime}-1)\omega,\omega,z$, then $[-\bar{l}_{k},\bar{l}_{k}]\cap([0,n-1]-m_{0}^{\prime}+1)$ is also NDR with respect to $x+(m_{0}^{\prime}-1)\omega,\omega,z$. Therefore, Lemma 5.10 guarantees that if $\Lambda(m^{\prime})=[-\bar{l}_{k},\bar{l}_{k}]$, then (6.2) also holds for any subintervals of $[-\bar{l}_{k},\bar{l}_{k}]$ with length exceeding than $\bar{l}_{k}/10$.
Consider $m\in[0,n-\bar{l}]$ such that $|m-m_{0}|\geq\bar{l}+2\exp{l^{2}}$. Then $[m,m+\bar{l}-1]\subset\{m^{\prime}:|m^{\prime}-m_{0}^{\prime}|\geq\exp(l^{2})\}$. Every point in $[m,m+\bar{l}-1]$ is contained in an interval of the form $\Lambda(m^{\prime})\cap[m,m+\bar{l}-1]$ for some $m^{\prime}\in[m-1,m+\bar{l}-1-l]$, on which the large deviations estimate applies. Note that we aimed to ensure that $m^{\prime}+[1,l]\subset[m,m+\bar{l}-1]$, since if a large deviations estimate holds on $m^{\prime}+[1,l]$, it does not necessarily hold on $m^{\prime}+[1,l]\subset[m,m+\bar{l}-1]$. On the other hand, we already noted that when a large deviations estimate holds on $m^{\prime}+[-\bar{l}_{k},\bar{l}_{k}]$ it also holds on $(m^{\prime}+[-\bar{l}_{k},\bar{l}_{k}])\cap[m,m+\bar{l}-1]$. Since $\bar{l}\geq(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}$, then
$|\Lambda(m^{\prime})|\leq\bar{l}_{(s-1)^{2}}\leq\bar{l}^{\nu/2}$
as required for the covering form of LDT. Therefore, (6.1) follows from (6.2) and the covering form of LDT. ∎
Lemma 6.4.
Let $\varepsilon\in(0,\frac{1}{5})$, $\nu$ as in LDT, and $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ as in Lemma 9.1. For any $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega_{0}\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$, $x_{0}\in\mathbb{T}^{d}\backslash\mathcal{B}_{n,\omega}$, and any eigenvalue $z_{0}=z_{j}^{[0,n-1]}(\omega_{0},x_{0})$, such that $L(\omega_{0},z_{0})>\gamma$, there exists an interval $I=I(\omega_{0},z_{0},x_{0})\subset[0,n-1]$,
$|I|<\exp((\log n)^{5\varepsilon}),$
such that for any $(\omega,x)\in\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d}$, $|\omega-\omega_{0}|,|x-x_{0}|<\exp(-\exp((\log n)^{2\varepsilon\nu}))$, we have
$\Big{|}u_{j}^{[0,n-1]}(\omega,x;s)\Big{|}<\exp(-\frac{\gamma}{4}\mathrm{dist}(s,I)),$
provided $\mathrm{dist}(s,I)>\exp((\log n)^{2\varepsilon\nu})$.
Proof.
Let $u$ be a normalized eigenvector corresponding to $z_{0}$. Let $m_{0}$ be such that
$|u(m_{0})|=\max\limits_{i\in(0,n-1)}|u(i)|.$
Then $m_{0}$ satisfies the assumptions of Lemma 6.3 because otherwise we can derive a contradiction based on the proof of Lemma 3.23. Let $\bar{l}=\lceil(\exp(\overline{C}_{(s-1)^{2}}(\log n)^{3\varepsilon\nu/2}))^{2/\nu}\rceil$. According to Lemma 6.3 (with $s$ instead of $m$), (6.1) holds for $s\in[0,n-1]\backslash I$, $I=[0,n-1]\cap[m_{0}-\bar{l}-2\exp(l^{2}),m_{0}+\bar{l}+2\exp(l^{2})]$, $l=\lceil(\log n)^{2\varepsilon}\rceil$. Then we can obtain the conclusion by applying Theorem 3.27 (with $\bar{l}$ instead of $l$). ∎
Lemma 6.5.
Let $\varepsilon\in(0,\frac{1}{5})$ and $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ be as in Lemma 6.4. For any $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega_{0}\in\mathbb{T}^{d}(p,q)\backslash\Omega_{n}$, $x_{0}\in\mathbb{T}^{d}(p,q)\backslash\mathcal{B}_{n,\omega}$, and any eigenvalue $z_{0}=z_{j}^{[0,n-1](\omega_{0},x_{0})}$, such that $L(\omega_{0},z_{0})>\gamma$, if we take $I=I(\omega_{0},z_{0},x_{0})$ as in Lemma 6.4, then
$\Big{|}z_{k}^{[0,n-1]}(\omega,x)-z_{j}^{[0,n-1]}(\omega,x)\Big{|}>\exp(-C|I|),$
for any $k\neq j$ and any $(\omega,x)\in\mathbb{T}^{d}(p,q)\times\mathbb{T}^{d}$, with
$|\omega-\omega_{0}|,\,\,|x-x_{0}|<\exp(-\exp((\log n)^{2\varepsilon\nu})).$
Proof.
According to the proof of Lemma 6.4, (6.1) holds for $s\in[0,n-1]\backslash I$, $I=[0,n-1]\cap[m_{0}-\bar{l}-2\exp(l^{2}),m_{0}+\bar{l}+2\exp(l^{2})]$, $l=\lceil(\log n)^{2\varepsilon}\rceil$. Then the conclusion follows from Lemma 3.28. ∎
In what follows, when we deal with the difference between two vectors of different dimensions, we make the dimension of a low dimensional vector equal to the dimension of a high dimensional vector by adding the 00 element. For example, $\xi=(x_{-5},x_{-4},\ldots,x_{-1},x_{0},x_{1},\ldots,x_{4},x_{5})$, $\upsilon=(y_{-7},y_{-6},\ldots,y_{-1},y_{0},y_{1},\ldots,y_{6},y_{7}),$ $n>m$, $\upsilon-\xi=(y_{-7},y_{-6},y_{-5}-x_{-5},\ldots,y_{-1}-x_{-1},y_{0}-x_{0},y_{1}-x_{1},\ldots,y_{5}-x_{5},y_{6},y_{7})$. When the columns of a matrix are greater than the rows of a vector, we deal with their product in the same way.
Lemma 6.6.
Let $\varepsilon\in(0,\frac{1}{5})$. With $\Omega_{n}$, $\mathcal{B}_{n,\omega}$ as in Lemma 6.4, applied on $[-(n-1),n-1]$ instead of $[0,n-1]$, let $\hat{\Omega}_{N_{0}}=\mathop{\cup}\limits_{n\geq N_{0}}\Omega_{n}$, ^$\hat{\mathcal{B}}_{N_{0},\omega}=\mathop{\cup}\limits_{n\geq N_{0}}\mathcal{B}_{n,\omega}$. For $n\geq N_{0}(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\hat{\Omega}_{N_{0}}$, $x\in\mathbb{T}^{d}\backslash\hat{\mathcal{B}}_{N_{0},\omega}$, if $L(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x))>\frac{3\gamma}{2}$ and $I=I(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x),x)$ then for any $n\leq n^{\prime}\leq\exp((\log n)^{\frac{1}{10\varepsilon}})$ there exists $j_{n^{\prime}}$, such that
$\Big{|}z_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x)-z_{j}^{[-(n-1),n-1]}(\omega,x)\Big{|}<\exp(-\frac{\gamma}{12}n),$ (6.3)
$\Big{\|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x)-u_{j}^{[-(n-1),n-1]}(\omega,x)\Big{\|}<\exp(-\frac{\gamma}{48}n).$ (6.4)
Furthermore, $I(\omega,z_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x),x)\subset[-\frac{3}{4}(n^{\prime}-1),\frac{3}{4}(n^{\prime}-1)]$ and
$\Big{|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x;s)\Big{|}<\exp(-\frac{\gamma}{17}\mathrm{dist}(s,I)),\,\,n\leq|s|\leq n^{\prime}.$
Proof.
According to Lemma 6.4, we get that
$\Big{|}u_{j}^{[-(n-1),n-1]}(\omega,x;\pm(n-1))\Big{|}<\exp(-\frac{\gamma}{4}(\frac{n-1}{2}-1)),$
$\Big{|}u_{j}^{[-(n-1),n-1]}(\omega,x;\pm(n-2))\Big{|}<\exp(-\frac{\gamma}{4}(\frac{n-1}{2}-2)),$
$\Big{|}u_{j}^{[-(n-1),n-1]}(\omega,x;\pm(n-3))\Big{|}<\exp(-\frac{\gamma}{4}(\frac{n-1}{2}-3)),$
and
$\Big{|}u_{j}^{[-(n-1),n-1]}(\omega,x;\pm(n-4))\Big{|}<\exp(-\frac{\gamma}{4}(\frac{n-1}{2}-4)).$
Then we obtain that
$\Big{\|}(\mathcal{E}_{[-(n^{\prime}-1),n^{\prime}-1]}^{\beta,\eta}(\omega,x)-z_{j}^{[-(n-1),n-1]}(\omega,x))u_{j}^{[-(n-1),n-1]}(\omega,x)\Big{\|}<8\exp(-\frac{\gamma}{4}(\frac{n-1}{2}-4)).$
According to Corollary 6.2 (with $\tilde{\varepsilon}=8\exp(-\frac{\gamma}{4}(\frac{n-1}{2}-4)))$, there exists $j_{n^{\prime}}$ such that $z_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x)\in\mathcal{D}(z_{j}^{[-(n-1),n-1]}(\omega,x);\tilde{\varepsilon}\sqrt{2})\cap\partial\mathbb{D}$. Thus,
$\Big{|}z_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x)-z_{j}^{[-(n-1),n-1]}(\omega,x)\Big{|}<2\tilde{\varepsilon}\sqrt{2}<\exp(-\frac{\gamma}{12}n).$
Applying Corollary 6.2 again (with $\hat{\varepsilon}=\exp(-\frac{\gamma}{12}n)$), one can obtain that
$\Big{\|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x)-u_{j}^{[-(n-1),n-1]}(\omega,x)\Big{\|}<\sqrt{2}\hat{\varepsilon}^{-1}\tilde{\varepsilon}<\exp(-\frac{\gamma}{48}n).$
From 6.4, we can get that
$\mathop{\sum}\limits_{|s|>n-1}\Big{|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x;s)\Big{|}^{2}<\Big{\|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x)-u_{j}^{[-(n-1),n-1]}(\omega,x)\Big{\|}^{2}<\exp(-\frac{\gamma}{24}n),$
which implies that $\mathop{\sum}\limits_{|s|\leq\frac{5(n-1)}{9}}\Big{|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x;s)\Big{|}^{2}>\frac{1}{2}.$
It follows from Lemma 6.4 that
$I(\omega,z_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x),x)\subset[-\frac{3(n-1)}{4},\frac{3(n-1)}{4}],$
and
$\Big{|}u_{j_{n^{\prime}}}^{[-(n^{\prime}-1),n^{\prime}-1]}(\omega,x;s)\Big{|}<\exp\Big{(}-\frac{\gamma}{4}\mathrm{dist}\big{(}s,[-\frac{3(n-1)}{4},\frac{3(n-1)}{4}]\big{)}\Big{)}<\exp(-\frac{\gamma}{17}\mathrm{dist}(s,I))$
for any $n-1\leq|s|\leq n^{\prime}-1$. ∎
Theorem 6.7.
Let $\varepsilon\in(0,\frac{1}{5})$ and $\hat{\Omega}_{N_{0}}$, ^$\hat{\mathcal{B}}_{N_{0},\omega}$ be as in Lemma 9.6 For any $N_{0}\geq C(\alpha_{i},p,q,\gamma,\varepsilon)$, $\omega\in\mathbb{T}^{d}(p,q)\backslash\hat{\Omega}_{N_{0}}$, $x\in\mathbb{T}^{d}\backslash\hat{\mathcal{B}}_{N_{0},\omega}$, the following statements hold. Let $n_{k}=n^{2^{k}}$. If $L(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x))>2\gamma$ and $I=I(\omega,z_{j}^{[-(n-1),n-1]}(\omega,x),x)\subset[-\frac{n-1}{2},\frac{n-1}{2}]$, then for each $k\geq 1$ there exist $j_{k}$ such that
$\Big{|}z_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)-z_{j}^{[-(n-1),n-1]}(\omega,x)\Big{|}<\exp(-\frac{\gamma}{16}n),$ (6.5)
$\Big{\|}u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)-u_{j}^{[-(n-1),n-1]}(\omega,x)\Big{\|}<\exp(-\frac{\gamma}{50}n),$
$\Big{|}u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x;s)\Big{|}<\exp(-\frac{\gamma}{60}\mathrm{dist}(s,I)),\,\,\frac{3(n-1)}{4}\leq|s|\leq n_{k}-1.$ (6.6)
Furthermore, for any $k^{\prime}\geq k\geq 0$,
$\Big{|}z_{j_{k^{\prime}}}^{[-(n_{k^{\prime}}-1),n_{k^{\prime}}-1]}(\omega,x)-z_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{|}<\exp(-\frac{\gamma}{16}n_{k}),$ (6.7)
$\Big{\|}u_{j_{k^{\prime}}}^{[-(n_{k^{\prime}}-1),n_{k^{\prime}}-1]}(\omega,x)-u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{\|}<\exp(-\frac{\gamma}{50}n_{k}).$
Particularly, the limits
$z(\omega,x):=\lim\limits_{k\rightarrow\infty}z_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x),\,\,u(\omega,x;s):=\lim\limits_{k\rightarrow\infty}u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x;s),\,\,s\in\mathbb{Z}$
exist and
$|u(\omega,x;s)|<\exp(-\frac{\gamma}{60}\mathrm{dist}(s,I)),\,\,|s|\geq\frac{3(n-1)}{4}.$ (6.8)
Proof.
According to Lemma 6.6, there exists $j_{k}$, $k\geq 0$, $j_{0}=j$, such that
$\Big{|}z_{j_{k+1}}^{[-(n_{k+1}-1),n_{k+1}-1]}(\omega,x)-z_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{|}<\exp(-\frac{\gamma}{12}n_{k}),$
$\Big{\|}u_{j_{k+1}}^{[-(n_{k+1}-1),n_{k+1}-1]}(\omega,x)-u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{\|}<\exp(-\frac{\gamma}{48}n_{k}),$
$L(\omega,z_{j_{k+1}}^{[-(n_{k+1}-1),n_{k+1}-1]}(\omega,x))>\frac{3\gamma}{2},$
$I_{k}=I(\omega,z_{j_{k+1}}^{[-(n_{k+1}-1),n_{k+1}-1]}(\omega,x),x)\subset[-\frac{3(n_{k}-1)}{4},\frac{3(n_{k}-1)}{4}],$
$\Big{|}u_{j_{k+1}}^{[-(n_{k+1}-1),n_{k+1}-1]}(\omega,x;s)\Big{|}<\exp(-\frac{\gamma}{17}\mathrm{dist}(s,I_{k}))<\exp(-\frac{\gamma}{18}\mathrm{dist}(s,I))$ (6.9)
for all $n_{k}-1\leq|s|\leq n_{k+1}-1$. Then for any $k^{\prime}\geq k\geq 0$, we have
$\Big{|}z_{j_{k^{\prime}}}^{[-(n_{k^{\prime}}-1),n_{k^{\prime}}-1]}(\omega,x)-z_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{|}<\sum_{i=k}^{k^{\prime}-1}\exp(-\frac{\gamma}{12}n_{i})<\exp(-\frac{\gamma}{16}n_{k})$
and
$\Big{\|}u_{j_{k^{\prime}}}^{[-(n_{k^{\prime}}-1),n_{k^{\prime}}-1]}(\omega,x)-u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{\|}<\sum_{i=k}^{k^{\prime}-1}\exp(-\frac{\gamma}{48}n_{i})<\exp(-\frac{\gamma}{50}n_{k}).$
This proves (6.5), (6.7), and the existence of the limits $z(\omega,x)$, $u(\omega,x)$.
From (6.9), we know that (6.6) holds for $n_{k-1}-1\leq|s|\leq n_{k}-1$. Then it is enough to prove that (6.6) holds for $\frac{3(n-1)}{4}\leq|s|\leq n_{k-1}$. For $21\leq\hat{k}\leq k-$ 2, due to
$\Big{\|}u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)-u_{j_{\hat{k}}}^{[-(n_{\hat{k}}-1),n_{\hat{k}}-1]}(\omega,x)\Big{\|}<\sum_{i=k}^{\hat{k}-1}\exp(-\frac{\gamma}{48}n_{i})<\exp(-\frac{\gamma}{50}n_{k}),$
we have
$\big{|}u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x;s)\big{|}$ $<\big{|}u_{j_{\hat{k}}}^{[-(n_{\hat{k}}-1),n_{\hat{k}}-1]}(\omega,x;s)\big{|}+\exp(-\frac{\gamma}{50}n_{k})$
$<\exp(-\frac{\gamma}{18}\mathrm{dist}(s,I))+\exp(-\frac{\gamma}{50}n_{k})$
$<\exp(-\frac{\gamma}{60}\mathrm{dist}(s,I)),$
for $n_{\hat{k}-1}-1\leq|s|\leq n_{\hat{k}}-1$. Due to $n_{k}=n^{2^{k}}$, $21\leq\hat{k}\leq k-$ 2, we have $n-1\leq n_{\hat{k}-1}-1\leq|s|\leq n_{\hat{k}}-1\leq n_{k}-1$. Thus, (6.6) holds for $n-1\leq|s|\leq n_{k}-1$. According to Lemma 6.4,
$\big{|}u_{j}^{[-(n-1),n-1]}(\omega,x;s)\big{|}<\exp(-\frac{\gamma}{17}\mathrm{dist}(s,I))\,\,\,\mathrm{for}\,\,\,\frac{3(n-1)}{4}\leq|s|\leq n-1.$
Therefore,
$\big{|}u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x;s)\big{|}$ $<\big{|}u_{j}^{[-(n-1),n-1]}(\omega,x;s)\big{|}+\exp(-\frac{\gamma}{50}n_{k})$
$<\exp(-\frac{\gamma}{17}\mathrm{dist}(s,I))+\exp(-\frac{\gamma}{50}n_{k})$
$<\exp(-\frac{\gamma}{60}\mathrm{dist}(s,I))$
for $\frac{3(n-1)}{4}\leq|s|\leq n-1$. In conclusion, (6.6) holds for $\frac{3(n-1)}{4}\leq|s|\leq n_{k}-1$.
Taking $k^{\prime}\rightarrow\infty$ in (6.7), one can obtain that
$\Big{\|}u(\omega,x)-u_{j_{k}}^{[-(n_{k}-1),n_{k}-1]}(\omega,x)\Big{\|}<\exp(-\frac{\gamma}{50}n_{k}).$
Similar to the proof of (6.6), (6.8) follows. ∎
7 Appendix
In order to make the structure of this paper clearer, we put some useful definitions and lemmas in this section.
7.1 Some Useful Notations
Definition 4.
[7, Definition 2.3] For any positive numbers $a,b$ the notation $a\lesssim b$ means $Ca\leq b$ for some constant $C>0$. By $a\ll b$ we mean that the constant $C$ is very large. If both $a\lesssim b$ and $a\gtrsim b$, then we write $a\asymp b$.
7.2 Avalanche Principle
Lemma 7.1.
[14, Proposition 2.2] Let $A_{1},\ldots,A_{m}$ be a sequence of arbitrary unimodular $2\times 2$-matrices. Suppose that
$\min_{1\leq j\leq m}\|A_{j}\|\geq\mu\geq m$ (7.1)
and
$\max_{1\leq j<m}\big{[}\log\|A_{j+1}\|+\log\|A_{j}\|-\log\|A_{j+1}A_{j}\|\big{]}<\frac{1}{2}\log\mu.$ (7.2)
Then
$\Big{|}\log\|A_{m}\cdots A_{1}\|+\sum_{j=2}^{m-1}\log\|A_{j}\|-\sum_{j=1}^{m-1}\log\|A_{j+1}A_{j}\|\Big{|}<C_{A}\frac{m}{\mu},$ (7.3)
where $C_{A}$ is an absolute constant.
7.3 Subharmonic Functions
Definition 5.
[13, Definition 2.3] Suppose $u:\Omega_{1}\times\cdots\Omega_{d}\rightarrow\mathbb{R}\cup\{-\infty\}$ is continuous. Then $u$ is said to be separately subharmonic, if for any $1\leq j\leq d$ and $z_{k}\in\Omega_{k}$ for $k\neq j$ the function
$z\rightarrow u(z_{1},\ldots,z_{k-1},z,z_{k+1},\ldots,z_{d})$
is subharmonic in $z\in\Omega_{j}$.
Lemma 7.2.
[13, Lemma 2.4] Let $u(x_{1},\ldots,x_{d})$ be a separately subharmonic function on $\mathbb{T}_{h}^{d}$. Suppose that $\sup\limits_{\mathbb{T}_{h}^{d}}u\leq M$. There are constants $C_{d}$, $C_{d,h}$ such that, if for some $10<\delta<$ 1 and some $L$
$\mathrm{mes}\{x\in\mathbb{T}^{d}:u<-L\}>\delta,$
then
$\sup\limits_{\mathbb{T}^{d}}u\leq C_{h,d}M-\frac{L}{C_{h,d}\log^{d}(C_{d}/\delta)}.$
7.4 Cartan’s Estimate
In what follows, we denote $\mathcal{D}(z_{0},r)=\{z\in\mathbb{C}:|z-z_{0}|<r\}$.
Definition 6.
[13, Definition 2.12] Let $H\geq 1$. For an arbitrary set $\mathcal{B}\subset\mathcal{D}(z_{0},1)\subset\mathbb{C}$ we say that $\mathcal{B}\in\mathrm{Car}_{1}(H,K)$ if $\mathcal{B}\subset\mathop{\cup}\limits^{j_{0}}_{j=1}\mathcal{D}(z_{j},r_{j})$ with $j_{0}\leq K$, and
$\sum_{j}r_{j}<e^{-H}.$ (7.4)
If $d\geq 2$ is an integer and $\mathcal{B}\subset\mathop{\prod}\limits_{j=1}^{d}\mathcal{D}(z_{j,0},1)\subset\mathbb{C}^{d}$, then we define inductively that $\mathcal{B}\in\mathrm{Car}_{d}(H,K)$ if for any $1\leq j\leq d$ there exists $\mathcal{B}_{j}\subset\mathcal{D}(z_{j,0},1)\subset\mathbb{C}$, $\mathcal{B}_{j}\in\mathrm{Car}_{1}(H,K)$ so that $\mathcal{B}^{(j)}_{z}\in\mathrm{Car}_{d-1}(H,K)$ for any $z\in\mathbb{C}\backslash\mathcal{B}_{j}$, here $\mathcal{B}^{(j)}_{z}=\{(z_{1},\ldots,z_{d})\in\mathcal{B}:z_{j}=z\}$.
Lemma 7.3.
[13, Lemma 2.15] Let $\varphi(z_{1},\ldots,z_{d})$ be an analytic function defined on a polydisk $\mathcal{P}=\mathop{\prod}\limits_{j=1}^{d}\mathcal{D}(z_{j,0},1)$, $z_{j,0}\in\mathbb{C}$. Let $M\geq\sup\limits_{z\in\mathcal{P}}\log|\varphi(z)|$, $m\leq\log|\varphi(z_{0})|$, $z_{0}=(z_{1,0},\ldots,z_{d,0})$. Given $H\gg 1$, there exists a set $\mathcal{B}\subset\mathcal{P}$, $\mathcal{B}\in\mathrm{Car}_{d}(H^{1/d},K)$, $K=C_{d}H(M-m)$, such that
$\log|\varphi(z)|>M-C_{d}H(M-m)$ (7.5)
for any $z\in\frac{1}{6}\mathcal{P}\backslash\mathcal{B}$. Furthermore, when $d=1$ we can take $K=C(M-m)$ and keep only the disks of $\mathcal{B}$ containing a zero of $\varphi$ in them.
$\mathbf{Acknowledgments}$
This work was supported in part by the NSFC (No. 11571327, 11971059).